### v3

Abstract
Many small to large organizations have adopted the Microservices Architecture (MSA) style to develop and deliver their core businesses. Despite the popularity of MSA in the software industry, there is a limited evidence-based and thorough understanding of the types of issues (e.g., errors, faults, failures, and bugs) that microservices system developers experience, the causes of the issues, and the solutions as potential fixing strategies to address the issues. To ameliorate this gap, we conducted a mixed-methods empirical study that collected data from 2,641 issues from the issue tracking systems of 15 open-source microservices systems on GitHub, 15 interviews, and an online survey completed by 150 practitioners from 42 countries across 6 continents. Our analysis led to comprehensive taxonomies for the issues, causes, and solutions. The findings of this study inform that Technical Debt, Continuous Integration and Delivery, Exception Handling, Service Execution and Communication, and Security are the most dominant issues in microservices systems. Furthermore, General Programming Errors, Missing Features and Artifacts, and Invalid Configuration and Communication are the main causes behind the issues. Finally, we found 177 types of solutions that can be applied to fix the identified issues. Based on our study results, we formulated future research directions that could help researchers and practitioners to engineer emergent and next-generation microservices systems.
keywords:
[…]
1 Introduction
The software industry has recently witnessed the growing popularity of the Microservices Architecture (MSA) style as a promising design approach to develop applications that consist of multiple small, manageable, and independently deployable services [1]. Software development organizations may have adopted or planned to use the MSA style for various reasons. Specifically, some of them want to increase the scalability of applications using the MSA style, while others use it to quickly release new products and services to the customers, whereas it is argued that the MSA style can also help build autonomous development teams [2, 3]. From an architectural perspective, a microservices system (a system that adopts the MSA style) entails a significant degree of complexity both at the design phase as well as at runtime configuration [4]. This implies that the MSA style brings unique challenges for software organizations, and many quality attributes may be (positively or negatively) influenced [5]. For example, service level security may be impacted because microservices are developed and deployed by various technologies (e.g., Docker containers [6]) and tools that are potentially vulnerable to security attacks [4, 7]. Data management is also influenced because each microservice needs to own its domain data and logic [8]. This can, for example, challenge achieving and managing data consistency across multiple microservices.
Zimmermann argues that MSA is not entirely new from Service-Oriented Architecture (SOA) (e.g., “microservices constitute one particular implementation approach to SOA – service development and deployment”) [9]. Similarly, Márquez and Astudillo discovered that some existing design rationale and patterns from SOA fit the context for MSA [10]. However, an important body of literature (e.g., [11, 3]) has concluded that there are overwhelming differences between microservices systems, monolithic systems, and traditional service-oriented systems in terms of design, implementation, testing, and deployment. Gupta and Palvankar indicated that even having SOA experience can lead to suboptimal decisions (e.g., excessive service calls) in microservices systems [12]. Hence, microservices systems may have an additional and specific set of issues. Borrowing the idea from [13], we define issues in this study as errors, faults, failures, and bugs that occur in a microservices system and consequently impact its quality and functionality. Hence, there is a need to leverage existing methods or derive new practices, techniques, and tools to address the specific and additional issues in microservices systems.
Recently, a number of studies have investigated particular issues (e.g., code smell [14], debugging [15], performance [16]) in microservices systems. Despite these efforts, there is no in-depth and comprehensive study on the nature of different types of issues that microservices developers face, the potential causes of these issues, and possible fixing strategies for these issues. Jamshidi et al. believed that this can be partially attributed to the fact that researchers have limited access to industry-scale microservices systems [3]. The empirical knowledge on the nature of issues occurring in microservices systems can be useful from the following perspectives: (i) understanding common issues in the design and implementation of microservices systems and how to avoid them, (ii) identifying trends in the types of issues that arise in microservices systems and how to address them effectively, (iii) experienced microservices developers can be allocated to address the most frequent and challenging issues, (iv) novice microservices developers can quickly be informed of empirically-justified issues and avoid common mistakes, and (v) the industry and academic communities can synergies theory and practice to develop tools and techniques for the frequently reported issues in microservices systems.
Motivating Example: We now contextualize the issues, causes, and solutions based on an example illustrated in Figure 1. The example is taken from the Spinnaker project, an open-source microservices project hosted on GitHub (see Table 1), and annotated with numbering to represent a sequence among the reported issue, its cause(s), and the solution(s) to resolve the issue. Figure 1 shows auxiliary information about the Spinnaker project, such as project description, stars, and contributors. As shown in the example, a contributor, typically a microservices developer, writes code and may provide additional details of the code through comments. Once the code is compiled, the contributor reports permission denied issue highlighting “pipeline save with Admin account fails with permission denied”. As the next step, the same or other contributor highlights the cause for such issue as “Spinnaker user does not have access to the service account”. As the last step, an individual or a community of developers provides a solution such as “Allow the admin users to save the accounts” to resolve the issue. We are only interested in analyzing issues that have been marked as closed to ensure that a solution to resolve the issue exists. As shown in Table 1, the Spinnaker project has 4,595 closed issues and 121 contributors.
Figure 1: An example of the issue, cause, and their solution
This work aims to systematically and comprehensively study and categorize the issues that developers face in developing microservices systems, the causes of the issues, and the solutions (if any). To this end, we conducted a mixed-methods empirical study following the guideline proposed by Easterbrook and his colleagues [17]. We collected the data by (i) mining 2,641 issues from the issue tracking systems of 15 open-source microservices systems on GitHub, (ii) conducting 15 interviews, and (iii) deploying an online survey completed by 150 practitioners to develop the taxonomies of the issues, their causes and solutions in microservices systems. The key findings of this study are:
  1. 1.
    The issue taxonomy consists of 19 categories, 54 subcategories, and 402 types, indicating the diversity of issues in microservices systems. The top three categories of issues are Technical Debt, Continuous Integration and Delivery, and Exception Handling.
  2. 2.
    The cause taxonomy consists of 8 categories, 26 subcategories, and 228 types, in which General Programming Errors, Missing Features and Artifacts, and Invalid Configuration and Communication are the most frequently reported causes.
  3. 3.
    The solution taxonomy consists of 8 categories, 32 subcategories, and 177 types of solutions, in which the top three categories of solutions for microservices issues are Fix Artifacts, Add Artifacts, and Modify Artifacts.
  4. 4.
    The overall survey findings confirm the taxonomies of the issues, their causes and solutions in microservices systems and also indicate no major statistically significant differences in practitioners’ perspectives on the developed taxonomies.
The empirically validated taxonomies of issues, causes, and solutions in this study may greatly benefit software developers. For instance, developers can gain valuable insights into the diverse range of issues that they may encounter. These taxonomies can also help developers understand the causes of issues and implement preventive measures, enabling them to make informed decisions about addressing the issues during the development process. Furthermore, the proposed taxonomies in this study can facilitate knowledge sharing, training, and ensure a shared understanding among developers about the issues, causes, and solutions in microservices systems.
This paper has extended our previous work [18] by adding two new research questions (RQ3 and RQ4) and expanding and enhancing the results of RQ1 and RQ2 with increased volume and variety of data and a mixed-methods approach. Specifically, we explored 10 more open-source microservices systems on GitHub (now 15 projects), interviewed 15 practitioners, and conducted an online survey with 150 microservices practitioners for getting their perspectives on the proposed taxonomies of the issues, their causes and solutions in microservices systems, as well as the mapping between the issues, causes, and solutions.
Our study makes the following key contributions:
  1. 1.
    We developed the taxonomies of the issues, their causes and solutions in microservices systems based on a qualitative and quantitative analysis of 2,641 issue discussions among developers on GitHub, 15 interviews, and an online survey completed by 150 practitioners.
  2. 2.
    We provided the mapping between the issues, causes, and solutions in microservices systems with promising research directions on microservices systems that require more attention.
  3. 3.
    We made the dataset of this study available online [19], which includes the data collection and analysis from GitHub and microservices practitioners, as well as detailed hierarchies of the taxonomies of issues, causes, and solutions, to enable replication of this study and conduct future research.
The remainder of the paper is structured as follows. Section 2 details the research methodology employed. Section 3 presents the results of our study. Section 4 discusses the relationship between the issues, causes, and solutions, along with the implications and the threats to the validity of our results. Section 6 reviews related work, and Section 7 draws conclusions and outlines avenues for future work.
2 Methodology
The research methodology of this study consists of three phases, as illustrated in Figure 2. Given the nature of this research and the formulated research questions (see Section 2.1) – issues, causes, and solutions in microservices projects, we decided to use a mixed-methods study. Our study collected data from microservices projects hosted on GitHub, interviews, and a web-based survey. During Phase 1, we derived the taxonomies of 386 types of issues, 217 types of causes, and 177 types of solutions by mining and analyzing microservices practitioners’ discussions in the issue tracking systems of 15 open-source microservices projects hosted on GitHub. During Phase 2, we interviewed 15 microservices practitioners to extend and verify the taxonomies and identified additional 14 types of issues, 20 types of causes, and 22 types of solutions. During Phase 3, we surveyed 150 microservices practitioners using a Web-based survey to validate the outcomes of Phase 1 and Phase 2, i.e., using practitioners’ perspectives and feedback to validate the extracted types of issues, their causes and solutions.
2.1 Research Questions
We formulated the following research questions (RQs).
RQ1: What issues do occur in the development of microservice systems?
Rationale: RQ1 aims to systematically identify and taxonomically classify the types of issues that occur in microservices systems. The answer to RQ1 provides a comprehensive understanding of the issues (e.g., the most frequent issues) of microservices systems.
RQ2: What are the causes of issues that occur in microservices systems?
Rationale: The aim of RQ2 is to investigate and classify the root causes behind the issues identified in RQ1 and map causes to issues. The answer to RQ2 helps practitioners avoid common issues in microservices systems.
RQ3: What solutions are proposed to fix issues that occur in microservices systems?
Rationale: The aim of RQ3 is to identify the solutions for the issues according to their causes and to develop the taxonomy of solutions. The answer to RQ3 helps to understand the fixing strategies for addressing microservices issues.
RQ4: What are the practitioners’ perspectives on the taxonomies of the identified issues, causes, and solutions in microservices systems?
Rationale: The taxonomies of issues, causes, and solutions developed in RQ1, RQ2, and RQ3 are based on 15 open-source microservices systems and interviewing 15 practitioners. RQ4 aims to evaluate the taxonomies of issues, causes, and solutions built in RQ1, RQ2, and RQ3 by conducting a relatively large-scale online survey.
Figure 2: An overview of the research method
2.2 Phase 1 - Mining Developer Discussions
This phase aims to systematically identify and synthesize the issues, their causes and solutions in open-source microservices systems on GitHub. For an objective and fine granular presentation of methodological details, this phase is divided into two steps, each elaborated below, based on the illustrative view in Figure 2.
2.2.1 Step A – Identify & Select MSA-based OSS Projects
The specified RQs require us to identify and select MSA-based OSS projects, representing a repository of developer discussions and knowledge, to extract the issues, their causes, and solutions. This means that the RQs guided the development of search strings based on the recommendations and steps from [20] for string composition to retrieve developer discussions on microservices projects deployed on GitHub [21]. We formulated the search string using the format [keyword-1 [OR logic] keyword-2 … [OR logic] … keyword-N], where keywords represented the synonyms as [‘micro service’ OR ‘micro-service’ OR ‘microservice’ OR ‘Micro service’ OR ‘Micro-service’ OR ‘Microservice’]. To extract MSA issues, we selected GitHub, which is one of the most popular and rapidly growing platforms for social coding and community-driven collaborative development of OSS. GitHub represents a modern genre of software forges that unifies traditional methods of development (e.g., version control, code hosting) with features of socio-collaborative development (e.g., issue tracking, pull requests) [22]. The variety and magnitude of OSS available on GitHub also inspired our choice to investigate the largest OSS platform in the world, with approximately 40 million users and 28 million publicly available project repositories.
Based on the search string, we searched for the title and description of the OSS projects deployed on the GHTorrent dump hosted on Google Cloud. The search helped us retrieve a total of 2,690 potentially relevant MSA-based OSS projects for investigation. To shortlist and eventually select the projects pertinent to the outlined RQs, we applied multi-criteria filtering [23], considering a multitude of aspects such as the popularity or perceived significance of a project in the developers’ community (represented as total stars), adoption by or interests of developers (total forks), and the total number of developers involved (total contributors) for the project. As shown in Figure 2 (Step B), we only selected the projects that have (i) more than 10 stars and forks, (ii) an English description, and (iii) three or more contributors. This led to a shortlist of 167 microservices projects. To eliminate the instances of potential false positives, i.e., avoiding bias in construct validity, such as misleading project names and mockup code, we contacted the top three contributors of each project via their emails to clarify about:
  1. 1.
    Correct Interpretation of the Project: Please confirm if our interpretation of your project (Project URL and Name as an identifier) is appropriate for its design and implementation based on MSA. Also, please help us clarify if this project (e.g., tool, framework, solution) supports the development of microservices systems or if this project is developed using MSA.
  2. 2.
    MSA-based Characteristics and/or Features of the Project: What features and/or characteristics of the project reflect MSA being used in the project?
  3. 3.
    Optional Question: Could you please help us identify (the names, URLs, etc. of) any other OSS projects that are designed or developed using MSA?
We contacted a total of 426 contributors, with 39 of them responded to our query. Based on their confirmation, we selected 15 MSA-based OSS projects (see Table 1). These 15 selected projects, based on the available data at the time of project selection, represent the highest number of issues, making them highly relevant to our study.
Table 1: Identified open-source microservices systems
Project Name #Issues #Contributors #Forks #Stars
Spinnaker 4595 121 1.2K 8.5K
Cortex 1120 226 681 4.7K
Jaeger 995 239 1.9K 15.8K
eShopOnContainers 986 157 8.9K 20.7K
Goa 930 97 485 4.7K
Light-4j 584 37 588 3.4K
Moleculer 473 102 497 5.1K
Microservices-demo 287 55 2.1K 3.1K
Cliquet 207 19 20 65
Deep-framework 174 12 75 537
Scalecube 130 20 90 547
Lelylan 123 7 93 1.5K
Open-loyalty 175 14 80 300
Spring PetClinic 69 32 1.4K 1.1K
Pitstop 39 15 490 890
2.2.2 Step B – Synthesize Issues, Causes, and Solutions
After the projects were identified, as illustrated in Figure 2, extracting and synthesising the issues was divided into the following five parts.
Raw Data Collection: We chose 15 microservices projects (see Table 1) as the source for building the dataset to answer our RQs. These 15 projects were chosen because they are significantly larger than other microservices projects hosted on GitHub. Hence, it is highly likely that their contributors had more discussions about the type of issues, causes, and solutions in issue tracking systems. The discussions relating to a software system can usually be captured in issue tracking systems [24]. We initially extracted 10,222 issue titles, issue links, issue opening and closing dates, and the number of contributors for each issue through our customized Python script (see the Raw Data sheet in [19]). We stored this information in MySQL and exported it into MS Excel sheets for further processing. We extracted only closed issues because it could increase the chances of answering all our RQs (e.g., solutions).
Issues Screening: The first author further scanned the 10,222 issues to check if an issue has been closed or is still open. All open issues were discarded because an open issue is ongoing with many of its causes unknown and most likely solution(s) not found. Furthermore, after selecting only the closed issues, the first author further eliminated (i) issues without a detailed description, (ii) general questions, opinions, feedback, and ideas, (iii) feature requests (e.g., enhancements, proposals), (iv) announcements (e.g., about new updates), (v) duplicated issues, (vi) issues that had only one participant, and (vii) stale issues. After this step, we got 5,115 issues (see the Selected Issues (Round 1) sheet in [19]). We had second round of screening on these 5,115 issues to check whether these issues are related to our RQs or not, and after comprehensively analyzing them we found 2,641 issues that were related to our RQs (see the Selected Issues (Round 2) sheet in [19]).
Pilot Data Extraction: To gain initial insights into the issues, two authors (i.e., the first and fourth) performed pilot data extraction based on 150 issues (5.67% of 2,641 issues). The authors focused on issue-specific data, such as issue description (i.e., textual details specified by contributors), type of issue (e.g., testing issue, deployment issue), and frequency of issue (i.e., number of occurrences). Pilot data extraction was checked by the second and third authors to verify and refine the details before final data extraction.
Data Extraction: Issues, causes, solutions, and their corresponding data were extracted based on the guidelines for mining software engineering data from GitHub [25]. We extracted issue titles, issue links, the number of participants for each issue, key points of issues, key points of causes, and key points of solutions (see the Raw Data and Initial Codes sheet in [19]). Data such as issue title, issue link, and number of participants for each issue were used to capture general information about developer discussions. On the other hand, key points of issues, causes, and solutions were used to document the data for answering RQ1, RQ2, RQ3.
Data Analysis: To synthesize the issues, we used the thematic analysis approach [26] to identify the categories of issues, causes, and solutions. The thematic analysis approach is composed of five steps. (i) Familiarizing with data: The first author repeatedly read the project’s contributor’s discussion and documented all discussed key points about issues, causes, and solutions. (ii) Generating initial codes: after data familiarization, the first author generated an initial list of codes from the extracted data (see the Initial Codes sheet in [19]). (iii) Searching for the types of issues, causes, and solutions: The first and second authors analyzed the initially generated codes and brought them under the specific types of issues,causes, and solutions. (iv) Reviewing types of issues, causes, and solutions: All the authors reviewed and refined the coding results with the corresponding types of issues, causes, and solutions. We separated, merged, and dropped several issues, causes, and solutions based on a mutual discussion between all the authors. (v) Defining and naming categories: We defined and further refined all the types of issues, causes, and solutions under precise and clear subcategories and categories (see Figure 4). Throughout this process, we engaged in the collaborative refinement and organization of issues, causes, and solutions into specific subcategories and categories. The naming and assignment of issues, causes, and solutions in microservices systems to the appropriate subcategories were conducted through consensus among all the authors, ensuring the development of a comprehensive and coherent taxonomy. Finally, we introduced three levels of categories for managing the identified issues, causes, and solutions. First, we organized the types of issues, causes, and solutions under a specific subcategory (e.g., service dependency in Service Design Debt). Then we arranged the subcategories under a specific category (e.g., Service Design Debt in Technical Debt).
To minimize any bias during data analysis, each step, i.e., classification, mapping, and documentation, were cross-checked and verified independent of the individual(s) involved in data synthesis. Documentation of the results as answers to RQs is presented as taxonomical classification of issues (Section 3.1), causes of issues (Section 3.2), solutions to address the issues (Section 3.3), and mapping of issues with their causes and solutions (Section 4.1), complemented by details of validity threats (Section 5).
2.3 Phase 2 - Conducting Practitioner Interviews
Since we aim to understand the issues, causes, and solutions of microservices systems from a practitioners’ perspective, we opted to conduct interviews to confirm and improve the developed taxonomies. The interview process consists of the following steps.
2.3.1 Preparing a Protocol
The first author conducted 15 online interviews with microservices practitioners through Zoom, Tencent Meeting, and Microsoft Teams. Before conducting actual interviews, we also conducted two pilot interviews with microservices practitioners to check the understandability and comprehensiveness of interview questions. However, we did not include their answers in our dataset. In total, we conducted 15 actual interviews, and each interview took 35-45 minutes. It is argued that conducting 12 to 15 interviews with homogeneous groups is enough to reach saturation [27]. In our case, after completing 15 interviews, we noticed a clear saturation in the responses to our interview questions, indicating that we had gathered enough information to address our interview objectives effectively. Therefore, we decided not to conduct any additional interviews. We conducted semi-structured interviews based on an interview guide (see the Interview Questionnaire sheet in [19]), which contains closed-ended questions related to participant demographics and open-ended questions related to general categories of microservices issues, causes, and solutions.
The interview process was comprised of three sections. In the first section, we asked 6 demographic questions to understand the interviewee’s background in microservices. We covered various aspects in this section, including the country of the practitioner, major responsibilities, overall experience in the IT industry, experience with implementing microservices systems, the work domain of the organization, and programming languages for developing microservices systems. In the second part, we asked three open-ended questions about the types of issues, causes of issues, and solutions to issues during the development of microservices systems. The purpose of this part was to allow the interviewees to spontaneously express their views about the issues developers face in developing microservices systems, the causes of the issues, and resolution strategies without the interviewer biasing their responses. In the third part, we presented the three taxonomies to the interviewees and asked them to indicate any missing issues, causes, and solutions that have not been explicitly mentioned. All three taxonomies come from identifying, analyzing, and synthesizing the developer discussions from 15 open-source microservices systems. The taxonomy of issues consists of 386 types of issues, 54 issue subcategories, and 18 issue categories. The taxonomy of causes consists of 217 types of causes, 26 cause subcategories, and 8 cause categories. The taxonomy of solutions consists of 171 types of solutions, 33 solution subcategories, and 8 solution categories. At the end of each interview, we thanked the interviewee and briefly informed them of our next plans.
2.3.2 Conducting Interviews
To recruit interviewees, we initially reached out to our professional contacts in each country via email. We also requested that they disseminate our interview request among their colleagues who they believed would be suitable candidates for the study. We informed the possible participants that this interview was entirely voluntary with no compensation. With this approach, we recruited 15 microservices practitioners from IT companies in 10 countries: Australia (1), Canada (3), China (1), Chile (1), India (1), Pakistan (2), Sweden (2), Norway (1), the United Kingdom (1), and the United States of America (2). We refer to the interviewees as P1 to P15. Most of the interviewees are mainly software architects and application developers. Their average experience in the IT industry is 10.33 years (Minimum: 5, Maximum: 16, Median: 10, Mode: 9, Standard Deviations: 3.26). The interviewees’ average experience in microservices is 5.33 years (Minimum: 3, Maximum: 8, Mode: 4, Median: 5, Standard Deviations: 1.67).
2.3.3 Data Analysis
We applied a thematic analysis method [28] to analyze the recorded interviews. Before applying the thematic analysis method, the first author prepared the text transcripts from audio recordings. The first author read the interview transcripts and coded them using the MAXQDA tool. We dropped several sentences unrelated to “microservices issues, causes, and solutions”. After removing the extraneous information from the transcribed interviews, the first author coded the interview transcripts’ contents to get the answers to the interview questions. To ensure the quality of the codes, the second author verified the initial codes created by the first author and provided suggestions for improvement. After the initial coding and subsequent enhancements, all the authors collaborated in analyzing the codes. Conflicts in the codes were addressed through collaborative discussions and brainstorming sessions. Finally, we generated a total of 28 types of issues (classified into 15 subcategories and 11 categories), 28 types of causes (classified into 15 subcategories and 7 categories), and 30 types of solutions (classified into 4 subcategories and 3 categories). During the interviews, we got 48 instances of issues, 31 instances of causes, and 36 instances of solutions. Among these instances, we identified 14 types of issues, 21 types of causes, and 23 types of solutions that were not part of the taxonomies we derived from the 15 open-source microservices systems. Later, we exported the analyzed interview data from the MAXQDA tool to an MS Excel sheet (i.e., the Interview Results sheet in [19]) to make the part of taxonomies of issues, causes, and solutions in microservices systems from the interview data.
Except for service size, operational and tooling overhead, and team management issues, all of the issues, causes, and solutions identified through the interviews can be classified under the existing categories (i.e., the output of Phase 1). The instances of issues mentioned by the interviewees include CI/CD Issues (9), Security Issues (6), Service Execution and Communication Issues (5), Database Issues (5), Organizational Issues (5), Testing Issues (5), Monitoring Issues (4), Performance Issues (4), Update and Installation Issues (3), Configuration Issues (1), and Technical Debt (1). The causes mentioned by the interviewees include Service Design and Implementation Anomalies (20), Poor Security Management (2), Legacy Versions, Compatibility, and Dependency Problems (2), Invalid Configuration and Communication Problems (2), General Programming Errors (2), Fragile Code (2), and Insufficient Resources (1). The solutions mentioned by the interviewees include Add Artifacts (32) and Upgrade Tools and Platforms (4).
Table 2: Overview of the interviewees and their demographic information
# Responsibilities Languages Domain Overall Exp. MSA Exp. Country
P1 Software Architect, Developer Python, Java, Node.JS E-commerce, Healthcare 14 Years 6 Years Sweden
P2 Developer Java with Spark E-commerce 9 Years 5 Years USA
P3 Software Architect Python, Go, Java E-commerce 13 Years 6 Years UK
P4 Software Architect, Developer Python, Java Educational ERP 7 Years 3 Years Pakistan
P5 Software Architect Python, Java Internet of Things 12 Years 7 Years Canada
P6 Software Architect, Developer Java, Node.JS, Python Healthcare 15 Years 8 Years Australia
P7 Software Engineer C#.Net E-commerce, Banking 10 Years 4 Years Canada
P8 DevOps Consultant Kotlin, Python Network applications 9 Years 6 Years Canada
P9 Software Architect Java Education, Healthcare 5 Years 4 Years Chile
P10 Application Developer, Architect Java Financial (Insurance) 10 Years 8 Years Sweden
P11 DevOps Consultant Java, Kotlin,Python Telecommunication 6 Years 3 Years Norway
P12 Application Developer Angular, Puthon Manufacturing 8 Years 4 Years China
P13 Principal Consultant Swift Transportation 9 Years 5 years USA
P14 Azure Technical Engineer JavaScript, Golang Embedded systems 12 Years 4 Years Pakistan
P15 Software Architect Ruby, UML Payment applications 16 Years 7 Years India
2.4 Phase 3 - Conducting a Survey
A questionnaire-based survey approach is used to evaluate the taxonomies of issues, causes, and solutions based on mining developer discussions and conducting practitioner interviews. We adopted Kitchenham and Pfleeger’s guidelines for conducting surveys [29] and used an anonymous survey to increase response rates [30].
2.4.1 Recruitment of Participants and Conducting the Pilot Survey
After the survey design, we needed to (i) select the survey participants and (ii) conduct a pilot survey for initial assessments (e.g., time taken, clarity of statements, and add, remove, and refine the questions). To select the potential respondents, we used the following contact channels to spread the survey broadly to a wide range of companies from various locations worldwide. The contact channels to recruit the potential participants included (i) professional contacts, researchers of industrial track publications, and authors of web blogs related to microservices, (ii) practitioners and their communities on social coding platforms (e.g., GitHub, Stack Overflow), and (iii) social and professional online networks (LinkedIn, Facebook, Twitter, Google Groups). In the survey invitation email, we also requested the potential participants to share the survey invitation with individuals or groups deemed as relevant participants. Before sending out the invitations, we ensured that we only contact individuals with experience in any aspects of MSA design, development, and/or engineering based on their professional profiles, such as code commits, industrial publications, and professional designations. Based on publicly available email IDs, first, we sent out survey invitations to only a selected set of 30 participants for a pilot survey. Out of the 30 participants contacted for the pilot survey, 10 replied (response rate 33.33%) from 7 countries. It should be noted that the results of the pilot survey were not incorporated into our primary survey data. However, the pilot survey helped us to refine the survey questionnaire in terms of restructuring the sections and rephrasing some questions for clarity of the survey to ensure that (i) the length of the survey is appropriate, (ii) the terms used in the survey questions are clear and understandable, and (iii) the answers to the survey questions are meaningful.
Figure 3: Overview of the demographics of survey participants
2.4.2 Conducting the Web-based Survey
We adopted a cross-sectional survey design, which is appropriate for collecting information at one given point in time across a sample population [29]. Surveys can be conducted in many ways, such as Web-based online questionnaires and phone surveys [31]. We decided to conduct a Web-based survey because these surveys can help to (i) minimize the time and cost, (ii) collect the responses from geographically distributed respondents, (iii) minimize time zone constraints, and (iv) save the effort of researchers to collect data in a textual, graphical, or structured format [31]. To document different types of responses while maintaining the granularity of information, we structured the questionnaire into a total of 12 questions organized under four sections (see the Survey Questionnaire sheet in [19]).
Demographics: We asked 6 demographic questions about the background information of the respondents to identify the (i) country or region, (ii) major responsibilities, (iii) overall work experience in the IT industry, (iv) work experience with microservices systems, (v) work domain of the organization, and (vi) programming languages and implementation technologies for developing microservices systems. The demographic information has been collected to (i) identify respondents who do not have sufficient knowledge about microservices, (ii) divide the results into different groups, and (iii) generalize the survey findings for the microservices research and practice community. We received a total of 156 responses, and we excluded 6 responses that were either randomly filled or filled by research students and professors who were not practitioners. In the end, we got 150 valid responses.
  • Countries: Respondents came from 42 countries of 6 continents (see Figure 3(a)) working in diverse teams and roles to develop microservices systems. The majority of them are from China (16 out of 150, 10.66%), Pakistan (13 out of 150, 8.66%), and India (9 out of 150, 6.00%).
  • Experience: We asked the participants about their experiences in the IT industry and the development of microservices systems. Figure 3(b) shows that the majority of the respondents (57 out of 150, 38.00%) have worked in the IT industry for more than 10 years and around one third of the respondents (44 out of 150, 29.33%) have worked with microservices systems for 1 to 3 years. We also received a considerable amount of responses in which practitioners have more than 10 years of experience working with microservices systems (20 out of 150, 13.33%).
  • Professional Roles: Figure 3(c) shows that the majority of the participants were application developer (62 out of 150, 41.33%), architect (40 out of 150, 26.66%), and DevOps engineer (29 out of 150, 19.8%). Note that one participant may have multiple major responsibilities in the company, and consequently, the sum of the percentages exceeds 100%.
  • Application Domains: Figure 3(d) shows the domains of the participants’ organizations where the microservice practitioners worked. Financial Systems (57 out of 150, 38.00%), E-commerce (29 out of 150, 19.33%), and Professional Services (29 out of 150, 19.33%) are the dominant domains. Note that one organization where a practitioner worked may have one or more application domains.
  • Programming Languages and Implementation Technologies: Figure 3(e) shows that 38 programming languages and technologies were used to develop microservices systems, in which “Java" (70 out of 150, 46.66%), “Python" (67 out of 150, 44.66%), and “GO" (39 out of 150, 26.00%) are the most frequently used languages for developing microservices systems.
Microservices Practitioners’ Perspective: To evaluate the taxonomies of issues, causes, and solutions, we asked six survey questions (both Likert scale and open-ended, see the Survey Questionnaire sheet in [19]) from microservices practitioners. We provided a list of 19 issue categories and asked survey participants to respond to each category on a 5-point Likert scale (Very Often, Often, Sometimes, Rarely, Never). Similarly, regarding causes, we provided 8 categories and asked practitioners to respond to each category on a 5-point Likert scale (Strongly Agree, Agree, Neutral, Disagree, Strongly Disagree). We also provided 8 categories of solutions and asked practitioners to respond to each category on a 5-point Likert scale. Along with these 5-point Likert scales, we added one option to know the familiarity of the survey respondents with the listed categories. We also asked three open-ended questions to identify the missing issue, causes, and solutions in the provided categories. All the open-ended responses were further analyzed through thematic classification and adjusted in the taxonomies.
2.4.3 Data Analysis
[…]
3 Results
This section presents the analyzed results of this study, addressing the four RQs outlined in Section 2.1. The analyzed results are further organized as categories (e.g., Technical Debt), subcategories (e.g., Code Debt), and types (e.g., code smell). We present categories in boldface, subcategories in italic, and types in small capitals. The relevant examples are provided as quoted messages along with their issue ID numbers to facilitate the traceability to our dataset (see the Initial Codes sheet in [19]). We report the types of issues in Section 3.1 (see Figure 4), the types of causes in Section 3.2 (see Table 3), the types of solutions in Section 3.3 (see Table 4), and practitioners’ perspective on these taxonomies in Section 3.4 (see Table 5).
3.1 Types of Issues (RQ1)
The taxonomy of issues in microservices systems is provided in Figure 4. The taxonomy of issues is derived by mining developer discussions (i.e., 2,641 instances of issues, see Section 2.2.2), conducting practitioner interviews (i.e., 48 instances of issues, see Section 2.3.3), and conducting a survey (9 instances of issues, see Section 3.4). Therefore, we got a total of 2,698 instances of issues. The results show that Technical Debt (687 out of 2698), Continues Integration and Delivery (313 out of 2698), and Service Execution and Communication (219 out of 2698) issues are most frequently discussed. It is worth noting that one issue may cover more than one topic. However, we consider the main topic of an issue where there is a high probability of identifying the causes or solutions. The number of issues in each issue type, subcategory, and category are also shown in Figure 4.
1. Technical Debt (687/2698, 25.46%): Technical Debt (TD) is “a metaphor reflecting technical compromises that can yield short-term benefit but may hurt the long-term health of a software system” [33]. This is the largest category in the taxonomy of microservices issues and includes a wide range of issues related to Code Debt and Service Design Debt. The interviewees also mentioned several issues regarding this category, and one representative quotation is depicted below.
[…]
The complexity introduces several types of technical debt both in the design and development phases of microservices systemsDeveloper, P2.
We identified and classified 19 types of TD issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Code Debt (658, 24.47%) refers to the issues of source code, which could adversely impact the code quality. This subcategory mainly gathers issues related to code refactoring, code smell, code formatting, excessive literals, and duplicate code. For instance, developers refactored the code of the Spinnaker project by “adding hal command for tweaking the component sizings, #11”. In addition, several other types of Code Debt, such as code formatting, excessive literals, and duplicate code also negatively affect the code legibility of microservices systems.
  • Service Design Debt (29, 1.07%) refers to the violation of adopting successful practices (e.g., MSA patterns) for designing open-source microservices systems. The issues in this subcategory are mainly related to service dependency, business logic issue, and design pattern issue. For instance, the developers of the moleculer project reported the issue of service design debt in which “Service A requires module A. If service A changed, the runner reloads, but if module A changed, the runner does not reload, #1873”.
2. Continuous Integration and Delivery (CI/CD) Issue (313/2698, 11.60%): CI/CD refers to the automation process that enables development teams to frequently develop, test, deploy, and modify software systems (e.g., microservices systems). Usually, a variety of tools and technologies are used to implement the CI/CD process. A wide range of CI/CD issues has been identified by mining microservices systems. However, we also found a few issues that the interviewees mentioned, and one representative quotation is depicted below.
[…]
The key issues of CI/CD for practitioners are many small independent code bases, multiple languages, frameworks, microservices integration, load testing, managing releases, and continued service updates”, DevOps Consultant, P8.
We identified and classified 55 types of CI/CD issues in 7 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Deployment and Delivery Issue (105, 3.89%) reports the problems that occur during the deployment and delivery of microservices systems. We identified 17 types of issues in this subcategory, which are mainly related to cd pipeline error, cd pipeline stage, halyard deployment, and deployment script errors. For example, regarding cd pipeline error, one contributor of the eShopOnContainers project highlighted that “Jenkins pipeline is failing with error 403, #990”.
  • Kubernetes Issue (74, 2.74%) reports the CI/CD issues specific to Kubernetes which is an open-source system for automatic deployment, scaling, and management of containerized applications (e.g., microservices systems). Most problems of the subcategory are related to general kubernetes, helm bake, kubernetes manifest errors. For example, the contributors of the eShopOnContainers project were “unable to list Kubernetes resources using default ASK to create a script, #688”).
  • Docker Issue (50, 1.85%): Docker is an open-source platform that helps practitioners with continuous testing, deployment, executing, and delivering applications (e.g., microservices systems). Our results indicate that most of the issues related to Dockers are docker image error, docker configuration error, and outdated container. For instance, the contributors of the eShopOnContainers project pointed out that the “building of the solution using docker-compose was failing, #2464” due to a docker configuration error.
  • Amazon Web Services (AWS) Issue (17, 0.63%): AWS provides an on-demand cloud computing platform for creating, testing, delivering, and managing applications. The most frequent types of this subcategory are general aws error and aws jenkins error. As an example, some developers of the Spinnaker project faced a situation in which “AWS Jenkins multi Debian package jobs fail to bake, #2429”.
  • Version Control Issue (16, 0.69%) is related to version control and management systems. In general, the major issues in this subcategory are related to Git, such as git plugin, master branch, and gitlab issues.
  • Google Cloud Issue (8, 0.29%): Google Cloud provides infrastructure services for creating and managing projects (e.g., microservices systems) and resources. This subcategory contains issues related to the Google Cloud platform for microservices systems. The most reported issues are gcp error gke error, and gce clone error. For instance, in the Spinnaker project, practitioners identified that “GCP: instance group port name Mapping is not working properly, #2496”.
  • Others (38, 1.40%): This subcategory gathers issues related to cloud driver error, virtual machine error, and spring boot error.
3. Exception Handling Issue (228/2698, 8.45%): Exception handling is used to respond the unexpected errors during the running state of software systems (e.g., microservices systems), and it helps to avoid the software system being crashed unexpectedly. This category represents the issues practitioners face when handling various kinds of exceptions in microservices systems. We identified and classified 44 types of Exception Handling issues in 5 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Unchecked Exception (81, 3.00%): These exceptions cannot be checked on the program’s compile time and throw the errors while executing the program’s instructions. We identified 10 types of issues in this subcategory, in which the top three types are null pointer exception, file not found exception, and runtime exception.
  • Checked Exception (77, 2.85%): These exceptions can be checked on the program’s compile time. Checked exceptions could be fully or partially checked exceptions. We identified 16 types of issues in this subcategory, mainly related to io exception, variables are not declared, and error handling.
  • Resource not Found Exception (37, 1.33%): These exceptions occur when some services cannot find the required resources for executing operations. We identified 8 types of issues in this subcategory. Most of them are related to attributes do not exist, no server group, and missing library.
  • Communication Exception (28, 1.03%): These are exceptions thrown when the client services cannot communicate with the producer services. We identified 7 types of issues in this subcategory, mainly related to http request exception and timeout exception.
  • Others (5, 0.18%): This subcategory gathers issues related to api exception, dependency exception, and thrift exception.
4. Service Execution and Communication Issue (219/2698, 8.11%): Communication problems are obvious when microservices communicate across multiple servers and hosts in a distributed environment. Services interact using e.g., HTTP, AMQP, and TCP protocols depending on the nature of services. The interviewees also mentioned a few issues regarding this category, and one representative quotation is depicted below.
[…]
The poor implementation of microservices communication is also a source of insecure communication, latency, lack of scalability, and errors and fault identification on runtime” (P7, Software Engineer).
We identified and classified 34 types of service execution and communication issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below
  • Service Communication (166, 6.15%): There are different ways of communication (e.g., synchronous communication, asynchronous message passing) between microservices. This subcategory covers the issues of service communication in which the majority of the issues are related to service discovery failure, http connection error, and grpc connection error.
  • Service Execution (27, 1.00%): This subcategory contains the issues regarding asynchronous communication, dynamic port binding, rabbitmq messaging, and service broker during service execution. These issues occur due to various reasons. For example, a dependency issue between microservices occurred when “integration commands were sent asynchronously, #1710”. We also found several issues regarding dynamic port binding. For instance, a server module of the light-4j project could not dynamically allocate a “port on the same host with a given range, #1742”.
  • Service Management (17, 0.63%): This subcategory covers the issues that occur in the distributed event store platform, service management platform, and service networking layer. The majority of the problems that happen in this subcategory are kafka bug, kafka josn format issue, and eks (elastic kubernetes service) error.
5. Security Issue (213/2698, 7.89%): Microservices provide public interfaces, use network-exposed APIs for communicating with other services, and are developed by using polyglot technologies and toolsets that may be insecure. This makes microservices a potential target for cyber-attacks; therefore, security in microservices systems demands serious attention. Mining microservices systems have identified a wide range of security issues. Among those issues, microservices practitioners also mentioned several other security issues during the interviews. One representative quotation is depicted below.
[…]
The other problem is securing microservices at different levels. Specifically, we deal with microservices-based IOT applications that have more insecure points than traditional ones” (P5, Solution Architect).
We identified and classified 37 types of security issues in 4 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Authentication and Authorization (123, 4.55%): Authentication is the process of identifying a user, whereas authorization determines the access rights of a specific user to system resources. We found that the majority of the authentication and authorization issues are related to handling authorization header (e.g., Basic Auth, OAuth, OAuth 2.0) and shared authentication issues. handling authorization header is generally used to implement authorization mechanisms. Our study found several issues related to Basic Auth, OAuth, and OAuth 2.0 header failure and the non-availability of these security headers. For example, the developers of the eShopOnContainers project reported the issue about “invalid_request on auth from Swagger for Location API, #1990”. Moreover, we found several issues regarding improper implementation of shared authentication methods in microservices systems.
  • Access Control (64, 2.37%) is a fundamental element in securing the infrastructure of microservices or any software systems. Access control could be role- or attribute-based in a microservices system. The major types of issues in this subcategory are managing credential setup and security policy violation.
  • Secure Certificate and Connection (43, 1.59%): Our study reports several issues regarding implementing security certificates and standards, such as SSL, TSL, and JWT, which are used to secure communication between client-server, service-to-service, or between microservices. For example, we found a jwt error and an ssl connection issue in the Goa and Spinnaker project respectively. We also found several other types of secure certificate and connection issues, such as security token expired, tls certificate issue, and expired certificate.
  • Encryption and Decryption (13, 0.48%) is used to convert plain text into ciphertext and ciphertext into plain text to secure the information. We identified three types of issues related to this subcategory are data encryption, data decryption, and configuration decryption.
6. Build Issue (210/2698, 7.78%): Build is a process of preparing an application program for software release by collecting and compiling all required source files. The outcome of this process could be several types of artifacts, such as binaries and executable programs. We identified and classified 20 types of build issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Build Error (141, 5.22%): We identified several types of build errors which can interrupt the build process of microservices systems. We found that the majority of build errors are related to build script, plugin compatibility, docker build fail, build pipeline error, build file server error, source file loading, and module resolution.
  • Broken and Missing Artifacts (59, 2.18%): These issues occur during the build process’s parsing stage when the build systems verify the required information (e.g., files, packages, designated locations) in the build script files before executing the build tasks. This subcategory mainly covers the issues related to missing properties, packages, and files, broken files, and missing objects. We also identified several other types of broken and missing artifacts which include missing ami (amazon machine image), missing base parameter in the client, and missing link attributes.
  • Others (10, 0.37%): This subcategory mainly gathers issues related to wrong use of universally unique identifier, and inconsistent data generated.
7. Configuration Issue (121/2698, 4.48%): Configuration is a process of controlling and tracking and making consistent all the required instances for software systems(e.g., microservices systems). Microservices systems have multiple instances and third party applications to configure. This category gathers configuration issues during microservices system development, implementation, and deployment phases. Besides identifying the configuration issues from the OSS microservices system, microservices practitioners also indicated configuration issues during the interviews. One representative quotation is depicted in the following.
[…]
The major challenge for me is the poor microservices’ configuration, which grows as the application size grows—the configuration effect on implementation and deployment phases of the microservices systems. The poor configuration of microservices may lead to increased latency and decrease the speed of microservices calls between different services” (P12, Application Developer).
We identified and classified 16 types of configuration issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly reported below.
  • Configuration Setting Error (61, 2.26%): This subcategory contains issues associated with the setup configuration of different types of servers, databases, and cloud infrastructure platforms. The main types of issues in this subcategory are server configuration error, database configuration error, and aks (azure kubernetes service) configuration error.
  • Configuration File Error (60, 2.22%): This subcategory covers the issues that mainly occur due to providing incorrect values in environment setting variables. The major types of issues in this subcategory are configuration mismatch, conflict in configuration file names, and incorrect file path.
8. Monitoring Issue (89/2698, 3.29%): The dynamic nature of microservices systems needs monitoring infrastructures to diagnose and report errors, faults, failure, and performance issues. This category reports issues related to monitoring microservices systems. Several interviewees also mentioned monitoring issues for microservices systems. One representative quotation is depicted in the following.
[…]
Microservices systems host containerized or virtualized across distributed private, public, hybrid, and multi-cloud environments. Monitoring highly distributed systems like microservices systems through traditional monitoring tools is a challenging experience because these tools only focus on a specific component or the overall operational health of the system” (P14, Azure Technical Engineer).
We identified and classified 17 types of monitoring issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Tracing and Logging Management Issue (60, 2.22%): One prominent challenge of monitoring microservices systems is the collection of logs from containers and distributed tracing. We identified 7 types of issues in this subcategory, mainly related to distributed tracing error, logging management error, and observability issue.
  • Health Check Issue (17, 0.63%): This subcategory deals with the problems related to the health monitoring of microservices systems. We identified six types of issues, mainly related to health check API error, health check fail, and health check port error.
  • Monitoring Tool Issue (12, 0.44%): We also identified several issue discussions, in which microservices practitioners discussed the problems of three monitoring tools, including zipkin issue, jenkins issue, and tcp/tt health check issue.
9. Compilation Issue (79/2698, 2.92%): This category reports compilation errors, which mainly occur when the compiler cannot compile source code due to errors in the code or errors with the compiler itself. We identified and classified 9 types of compilation issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Illogically Symbols (60/2698, 2.22%): These issues occur when developers use illegal characters or incorrect syntax during the coding, for instance, syntax error, invalid parent id, and unexpected end of file.
  • Wrong Method Call (19/2698, 0.70%): These issues occur when the compiler tries to search for definitions of methods by invoking them through method calls and finds wrong parameter, wrong method call, and incorrect values.
10. Testing Issue (77/2698, 2.85%): Microservices systems pose significant challenges for testing because of many services, inter-communication processes, dependencies, network communication, and other factors. Among those issues, microservices practitioners also mentioned several other testing issues during the interviews. One representative quotation is depicted in the following.
[…]
Testing is another issue that I think is more challenging in microservices systems. I also think deploying each microservice as a singular entity and testing them is tedious and brings several problems. For example, testing coordination among multiple microservices when deploying one service as a singular entity” (P2, Application Developer).
We identified and classified 20 types of testing issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Test Case Issue (45, 1,66%): This subcategory covers the problems with test cases written to evaluate the expected output compliance with specific requirements for the microservices systems. Most of them are related to faulty test case, missing test case, and syntax error in test case.
  • Code and Component Test (21, 0.77%): This subcategory deals with the issues mainly related to debugging, api testing, and missing design test of microservices systems.
  • Application Test (11, 0.40%): This subcategory gathers the issues related to overall microservices application testing, which include load test case, broken integration test, and application security testing issues.
11. Documentation Issue (75/2698, 2.77%): Documentation for microservices systems may suffer from several problems. We identified and classified 8 types of documentation issues in two subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Insufficient Document (49, 1.81%): This subcategory covers the problems related to outdated document, broken images, and inappropriate examples. For instance, one contributor of the eShopOnContainers project mentioned the issue of “Out of date wiki guide for vs2015, #2030”.
  • Readability Issue (26, 0.95%): This subcategory is related to readability problems with provided documentation. The leading types of readability issues are poor readability, missing readme file, and old readme file.
12. Graphical User Interface (GUI) Issue (70/2698, 2.59%): This category reports the problems that can wreck the GUI of microservices systems. We identified and classified 68 types of GUI issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Broken User Interface Elements (38, 1.41%) are dysfunctional User Interface (UI) elements (e.g., buttons or text fields) that can become the reason for inconsistencies in the page layout across different devices (e.g., mobile and desktop browsers). This subcategory represents the faults mainly related to front end crash, uneditable contents, and broken images in ui.
  • Missing Information and Legacy UI Artifacts (32, 1.19%): This subcategory contains the problems of wrong and incomplete information along with outdated UI artifacts, mainly related to wrong gui display, displaying incomplete information, and selection not working.
13. Update and Installation Issue (68/2698, 2.52%): This category gathers the identified problems related to update and installation of the packages, libraries, tools, and containerization platforms required to develop and manage microservices systems. We identified and classified 16 types of update and installation issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Update Error (45, 1.67%): This subcategory represents the errors in which developers face the issues of outdated packages, platforms, technologies, and backward compatibility. The leading types of issues are outdated install packages ,backward compatibility issue, and json update error.
  • Installation Error (23, 0.85%): The development of microservices systems can be interrupted because of failure to install required languages, packages, and platforms. The top three types of issues are language package installation error, npm (node package manager) error, and gke (google kubernetes engine) installation error.
14. Database Issue (65/2698, 2.40%): The ownership of the microservices system database is usually distributed, and most of the microservices are autonomous and have a private data store relevant to their functionality. The distributed nature of microservices systems brings challenges like database implementation, data accessibility, and database connectivity. The microservices practitioners also mentioned several other database issues during the interviews. One representative quotation is depicted below.
[…]
Relational database for microservices systems. This issue occurs when we migrate from monolithic applications to microservices systems. It was mainly because of the missing transaction management system for getting data from the database of the old application (that was a relational database) through the microservices application” (P6, Software Architect, Developer).
We identified and classified 24 types of database issues in 3 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below
  • Database Connectivity (26, 0.96%): This subcategory refers to the issues that occur while establishing a database connection. The leading types of issues in this subcategory are sql container failure , sql transient connection failure , and database creation failure.
  • Database Query (31, 1.14%): The errors in this subcategory cover search query issues during the development of microservices systems. Such issues may cause a database performance bottleneck. The leading types of issues in this subcategory are wrong query, elasticsearch database error, and database search error.
  • Others (8, 0.29%): Other types of database issues that cannot be classified into the above subcategories are included in this subcategory, which are mainly related to database migration, database storage, and database adapter. textcolorredFor example, one developer of the Spinnaker project mentioned that “DB adapter issue - can not perform a textual search with a list, #1374”.
15. Storage Issue (54/2698, 2.00%): This category reports storage space problems during the development, execution, and management of microservices systems. We identified and classified 13 types of storage issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]), and each subcategory is further described below.
  • Storage Size Constraints (48, 1.78%): Different microservices have different data storage requirements. This subcategory covers the storage size constraints, mainly related to lack of main memory, cache issue, and storage backend failure.
  • Large Data Size (6, 0.21%): This subcategory gathers the issues related to large data size, including large image size, large message size, and large file.
16. Performance Issue (45/2698, 1.67%): Microservices systems offer various advantages over monolithic systems. However, several types of issues wreak the performance of microservices systems. The interviewees also mentioned performance overhead as an issue, and one representative quotation is depicted below.
[…]
Unlike a monolithic application whose deployment and management are seemingly easier due to centralized control and monitoring, a microservices-based application has numerous independent services that may be deployed on different infrastructures and platforms. Such an aspect increases its performance overhead. I also think that microservices systems consume more resources, creating a heavy burden for servers. In the response achieving the performance goal become questionable” (P15, Software Architect).
We identified and classified 16 types of performance issues in three subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Service Response Delay (28, 1.03%): This subcategory gathers the types of issues regarding delay in service response that ruins the performance of microservices systems, such as long wait, inconsistent payloads, and slow query.
  • Resource Utilisation (13, 0.48%): In this subcategory, we collected the issues that degrade the performance of microservices systems due to resource utilisation. These issues are mainly related to load balancer error, high cpu usage, and rate limiting error. One example of the load balancer issue mentioned by a Spinnaker project contributor is “Failed to create a Load Balancer when creating a new application, #1583”.
  • Lack of Scalability (4, 0.14%): Scalability is the system’s ability to respond the user demands and change workload by adding or removing resources. This subcategory gathers the issues related to scalability that can hinder microservices system growth, for instance, scale to cluster error and circuit breaker issue.
17. Networking Issue (41/2698, 1.51%): Deploying, executing, and communicating in microservices systems over the network is complex. It is observed that many problems may disrupt the network. We identified and classified 20 types of networking issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Hosting and Protocols (22, 0.81%): This subcategory represents the issues related to hosting protocols, ports, and topologies for microservices systems. The leading types of issues in this subcategory are localhost error, ip address issue, and udp (user datagram protocol) discovery error.
  • Service Accessibility (19, 0.70%): This subcategory of issues represent the cases where microservices practitioners face service accessibility problems. The leading types of issues are webhook error, broken urls, and dns (domain name system) error.
18. Typecasting Issue (35/2698, 1.29%): This category is related to the typecasting issues that occur when assigning a value of one primitive data type to another type. We identified and classified 9 types of typecasting issues in 2 subcategories (see Figure 4 and the Issue Taxonomy sheet in [19]). Each of them is briefly described below.
  • Type Conversion (20, 0.74%): This subcategory deals with the issues when variables are not correctly converted from one type to another. The top three types of conversion issues are identity conversion, boxing conversion, and enumeration validation issue.
  • Narrow/Wide Conversion (15, 0.55%): These issues occur when the compiler converts variables of a larger type into a smaller type (e.g., double to float) or a smaller type into a larger type (e.g., float to double). The top two types of narrow/wide conversion issues are narrowing primitive conversion and narrowing reference conversion.
19. Organizational Issue (7/2698, 0.25%): We derived this category based on the interviewees’ feedback on the taxonomy of the issues (see Figure 4 and the Issue Taxonomy sheet in [19]). The interviewees only mentioned three types of issues in this category, including team management, operational and tooling overhead, and service size. We depicted two representative quotations below.
[…]
One of the critical challenges in organizations is team management according to available people, their expertise, and their working habits” (P1, Software Architect, Developer).
[…]
Creating a reasonable size for each microservices’ (I mean, each microservice should have sufficient responsibilities). This is a bit tricky because most of the issues are rooted here” (P6, Software Architect, Developer).
Key Findings of RQ1: We identified 2,641 instances of issues by mining developer discussions in 15 open-source microservices systems with 48 instances of issues mentioned by the interviewees and 9 instances of issues mentioned by the survey participants, which are 2,698 issues in total. The issue taxonomy consists of 19 categories, 54 subcategories, and 402 types, indicating the diversity of the issues in microservices systems. The majority of issues are related to Technical Debt (25.46%), CI/CD (11.60%), and Exception Handling (8.45%).
3.2 Causes of Issues (RQ2)
The taxonomy of causes of microservices issues is provided in Table 3. It is worth mentioning that not all the issue discussions provide the information about their causes. Therefore, we identified 2,225 issue discussions containing information about the causes. The taxonomy of causes is derived by mining developer discussions (i.e., 2,225 instances of causes), conducting practitioner interviews (i.e., 31 instances of causes, see Section 2.3.3), and conducting a survey (i.e., 11 instances of causes, see Section 3.4). Hence, we got a total of 2,267 instances of causes. We identified a total of 228 types of causes that can be classified into 8 categories and 26 subcategories. Due to space limitations, we only list the top two types of causes for each subcategory in Table 3. The detail of the types of causes can be found in the dataset [19]. The results show that General Programming Error (860 out of 2267), Missing Features and Artifacts (386 out of 2267), and Invalid Configuration & Communication Problems (382 out of 2267) are the top three categories of causes. Each cause category is briefly discussed below.
1. General Programming Error (GPE) (860/2267, 37.93%): This category captures the causes that are based on a broad range of errors occurred in different phases of microservices system development, such as coding (e.g., syntax errors), testing (e.g., incorrect test cases), and maintenance (e.g., wrong examples in documentation) phases. We identified and classified 78 types of GPE causes in six subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
  • Compile Time Error (377, 16.62%): This subcategory gathers the causes of issues in which microservices practitioners violate the rules of writing syntax for microservices systems codes. These causes must be addressed before the program can be compiled. We identified 16 types of causes in this subcategory in which the top three are semantic error, syntax error in code, and variable mutations.
  • Erroneous Method Definition and Execution (262, 11.55%): The causes in this subcategory are related to incorrect or partly correct definitions and executions of methods associated with object messages. Generally, methods are referred as class building blocks linked together for sharing and processing data to produce the desired results. We identified 27 types of causes in this subcategory in which the top three are lack of cohesion in methods, long message chain, and wrong parameterization.
  • Incorrect Naming and Data Type (157, 6.92%): This subcategory covers the causes related to choosing incorrect names and data types for identifiers, methods, packages, and other entities in the source code. Our taxonomy contains 18 types of causes in this subcategory, and among them wrong data conversion, wrong data type, and wrong use of data types are the top three types of causes.
  • Testing Error (25, 1.10%): This subcategory covers the causes behind testing issues in microservices systems. In this subcategory, we identified 6 types of causes in which the top two types of causes are incorrect test case and incorrect syntax in test cases.
  • Poor Documentation (22, 0.97%): The documentation of software systems may contain critical information that describes the software product capabilities for system stakeholders. We identified 7 types of causes in this subcategory, in which the top three are typo in documents, and wrong example in documents.
Figure 4: A taxonomy of issues in microservices systems
  • Query and Database Issue (11, 0.48%): This subcategory contains the causes behind database issues in microservices systems. We identified 4 types of causes in this subcategory, and the top three types of causes are wrong query parameters, missing query parameters, and incorrect querying range.
2. Missing Features and Artifacts (MFA) (386/2267, 17.02%): This category represents the causes behind microservices systems issues that occur due to missing required features, packages, files, variables, and documentation and tool support. We identified and classified 27 types of MFA causes in four subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
  • Missing Features (186, 8.20%) denotes the nonexistence of system functionality in microservices systems. In this subcategory, we identified 14 types of causes, in which the top two types of causes are missing required system features and missing security features.
  • Missing Documentation and Tool Support (104,4.58%): Proper documentation and tool support is vital to keep the record of and track changes between system requirements, architecture, and source code. It also guide various system stakeholders (e.g., architects, developers, end-users) regarding design, architecture, and coding standards use in microservices systems. Moreover, several types of tool support are also necessary for different phases of microservices system development (e.g., development and deployment). The absence of proper documentation and tool support can bring several types of issues to microservices systems. This subcategory contains 4 types of missing documentation and tool support causes, and mong them missing readme file and missing development and deployment tool support are identified as the leading causes.
  • Missing Packages and Files (68, 2.99%): This subcategory groups the causes related to absence of required resources, packages, and files for developing, deploying, and executing microservices systems. We collected 9 types of missing packages and files causes, in which the leading three types of causes are missing resource, missing required package, and missing api.
  • Missing Variables (28, 1.23%): A few missing variables are also identified as the causes for several microservices issues. Compilers throw the error messages of missing variables if variables are set to a nonexistent directory or have the wrong names. This subcategory contains 7 types of missing variables causes, and among them missing environment variable and missing properties are identified as the leading causes.
3. Invalid Configuration and Communication (ICC) Problem (382/2267, 16.85%): Considering a large number of microservices, their distributed nature, and third-party plugins, microservices systems need to be configured for complete business operations properly. Each microservice has its instances and process, and services interact with each other using several inter-service communication protocols (e.g., HTTP, gRPC, message brokers AMQP). One of the interviewees also mentioned the following cause regarding invalid configuration and communication.
[…]
Microservices systems typically use one or more infrastructure and 3rd party services. Examples of infrastructure services include a service registry, a message broker, and a database server. During the configuration of microservices, a service must be provided with configuration data that tells it how to connect to the external or 3rd party services — for example, the database network location and credentials” (P1, Software Architect, Developer).
We identified and classified 29 types of causes in two subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
  • Incorrect Configuration (290, 12.79%): Configuration management in microservices systems is a hefty task because microservices are scattered across multiple servers, containers, databases, and storage units. Each microservices may have multiple instances. Therefore, an incorrect configuration may lead to several types of errors. This subcategory contains 16 causes for different types of microservices issues, in which the leading two causes are incorrect configuration setting and wrong connection closure.
  • Server and Access Problem (92, 4.05%): Each microservices acts as a miniature application that communicates with each other. We need to configure the infrastructure layers of the microservice system for sharing different types of resources. A poor configuration may lead to problems of accessibility for servers and other resources that bring multiple issues in microservices systems. In this subcategory, we identified 13 types of causes, in which the leading three types of causes are transient failure, service registry error, and wrong communication protocol.
4. Legacy Versions, Compatibility, and Dependency (LC&D) Problem (222/2267, 9.79%): This category represents a broad range of causes arising from outdated repositories, applications, documentation versions, development and deployment platforms, APIs, libraries, and packages. One of the interviewees also mentioned several causes regarding microservices issues, especially compatibility and dependency. One representative quotation is depicted below.
[…]
Along with the legacy code version, some of the critical reasons for the microservices issues are i) no clear strategy for code repository and branching, the mix of technologies (each team uses its way of development), dependencies on other services which are not released yet but going for integration and load testing, development pace varies from team to team, lack of centralized release manager, incompatibility of a new version of the service with previous services.” (P8, DevOps Consultant).
We identified and classified 28 types of causes in five subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
  • Compatibility and Dependency (59, 2.60%): A typical microservices system consists of several independent services running on multiple servers or hosts [34]. However, some microservices also depend on other microservices to complete business operations. Usually, practitioners ensure the compatibility (e.g., backward compatibility) of each microservice with previous versions of the microservices systems during the upgrading. In this subcategory, we identified 5 types of causes, in which the leading three types of causes are compatibility error, and outdated dependency.
  • Outdated and Inconsistent Repositories (53, 2.33%): This subcategory gathers the causes which are the source of the issues that occur when the online code repository of version control systems has updated the local files repository. The most frequent causes in this subcategory are old repository version, old dev branch, and version conflicts.
  • Outdated Application and Documentation Version (45, 1.98%): We identified several causes arising from using outdated applications and documentation versions in the selected systems, which are mainly related to document not updated, and deprecated software version.
  • Outdated Development and Deployment Platforms (38, 1.67%): Development and deployment platforms help developers build, test, and deploy microservices systems efficiently. We identified 10 types of causes in this subcategory, and among them old Kubernetes version and old visual studio version are the top two leading type of causes.
  • Outdated APIs, Libraries, and Packages (27, 1.09%): This subcategory covers the causes arising from the usage of outdated APIs, libraries, and packages. The most frequent types of causes in this subcategory are outdated version of library and old version of package.
5. Service Design and Implementation Anomalies (SD&IA) (174/2267, 7.67% ): This category covers the causes of issues when microservices practitioners cannot address the associated complexity of distributed systems at the design level. The interviewees also mentioned several causes related to service design and implementation anomalies, and one representative quotation is depicted below.
[…]
According to my experience, the primary reason behind the design issues is the distributed nature of microservices systems, which is becoming increasingly complicated with growing systems—especially where multiple subsystems also need to be integrated” (P7, Software Engineer).
We identified and classified 17 types of SD&IA causes in 2 subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
  • Code Design Anomaly (130, 5.73%): Code design anomalies refer to poorly written code that may lead to several problems (e.g., difficulties in maintenance or future enhancements) in microservices systems. Our taxonomy gets 12 types of causes in this subcategory and among them poor code readability, messy code, and data clumps are the top three types of causes.
  • System Design Anomaly (44, 1.94%): System design anomalies refer to poorly designed microservices architecture that may lead to maintainability, scalability, and performance issues. This subcategory covers 5 types of causes. Among them, the top three types of causes are wrong dependencies chain, lack of asrs, and wrong application decomposition.
6. Poor Security Management (PSM) (126/2267, 5.55%): Microservices systems are distributed over data centres, cloud providers, and host machines. The security of microservices systems is a multi-faceted problem that requires a layered solution to cope with various types of vulnerabilities. The interview participants also mentioned several causes for security issues, and one representative quotation is depicted below.
[…]
I think the basic reasons behind the security issues are poor understanding of microservices architecture, large attack surfaces (many distributed points), error-prone encryption techniques while services are communicating, and insecure physical devices” (P5, Solution Architect).
We identified and classified 21 types of PSM causes in 3 subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
  • Coding Level (55, 2.42%): This subcategory collects the causes where strict security principles and practices are not followed to prevent potential vulnerabilities while writing code of microservices systems. This subcategory covers 7 types of coding level causes, and among them, the top three types of causes are unsafe code, malformed input, and wrong implementation of security api.
  • Communication Level (40, 1.76%): Given the polyglot and distributed nature of microservices systems, practitioners need to secure the inter-microservices communication. This subcategory covers 10 types of communication level causes, and among them, the top two types of causes are security dependencies and wrong access control.
  • Application Level (29, 1.27%): The application level of security refers to security practices implemented at the interface between an application and various components (e.g., databases, containers) of microservices systems. This subcategory only contains 3 types of causes that are insecure configuration management, missing security features, and violation of the security policies.
7. Insufficient Resources (IR) (93/2267, 4.10% ): Microservices systems are at risk of delivering the required outcome without sufficient resources. We identified and classified 4 types of IR causes in two subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
  • Memory Issue (68, 2.99%): Microservices systems are developed by using multiple languages (e.g., Java, Python, C++) and platforms (e.g., containers, virtual machines). Some languages and platforms consume more memory than others. For instance, C/C++ consumes less memory than Java, and Python and Perl consume less memory than C/C++. This subcategory covers 2 types of causes that are limited memory for process execution, and ide problem with memory.
  • Lack of Human Resources, Tools and Platforms (25, 1.10%): This subcategory deals with the causes related to tools and platforms support for microservices systems. Only 3 types of causes related to this subcategory are identified, including lack of tool support, anddeployment platform problem.
8. Fragile Code (FC) (24/2267, 1.05%): Fragile code refers to code that is difficult to change, and a minor modification in fragile code may break the service or module. We identified and classified 6 types of FC causes in two subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
  • Poor Implementation of Code (20, 0.88%): This subcategory gathers the causes of poor code quality, which exhibits the buggy behaviour of microservices systems. We identified 3 types of causes in this subcategory, including poor object-oriented design, poor code reusability, and unnecessary code.
  • Poor Code Flexibility (11, 0.48%): Code flexibility is important for long-lived microservices project code bases, however, we found a few issues that occurred due to poor code flexibility. The 3 types of causes in this subcategory are divergent changes, delayed refactoring, and poorly organized code.
Table 3: Taxonomy of causes of issues in microservices systems
Category of Causes Subcategory of Causes Type of Causes
General Programming Error (GPE) (860) Compile Time Error (377) Semantic Error (205)
Syntax Error in Code (112)
Erroneous Method Definition and Execution (262) Lack of Cohesion in Methods (38)
Long Message Chain (36)
Incorrect Naming and Data Type (157) Wrong Data Conversion (39)
Wrong Data Type (20)
Testing Error (25) Incorrect Test Case (20)
Wrong Implementation of Security Pattern (5)
Poor Documentation (22) Typo in Documents (9)
Wrong Example in Documents (7)
Query and Database Issue (13) Wrong Query Parameters (6)
Missing Query Parameters (3)
Missing Features and Artifacts (MFA) (386) Missing Features (186) Missing Required System Features (111)
Missing Security Features (37)
Missing Documentation and Tool Support (104) Missing Readme File (88)
Missing Tool Support (14)
Missing Packages and Files (68) Missing Resource (41)
Missing Required Package (15)
Missing Variables (28) Missing Environment Variable (14)
Missing Properties (8)
Invalid Configuration and Communication (ICC) Problem (382) Incorrect Configuration (285) Incorrect Configuration Setting (227)
Wrong Connection Closure (24)
Server and Access Problem (92) Transient Failure (21)
Service Registry Error (24)
Legacy Versions, Compatibility, and Dependency (LC&D) Problem (222) Compatibility and Dependency (59) Compatibility Error (42)
Outdated Dependency (9)
Outdated and Inconsistent Repositories (53) Old Repository Version (38)
Old DEV Branch (7)
Outdated Application and Documentation Version (45) Document not Updated (33)
Deprecated Software Version (6)
Outdated Development and Deployment Platforms (38) Old Kubernetes Version (7)
Old IDE Version (6)
Outdated APIs, Libraries, and Packages (27) Outdated Version of Library (15)
Old Version of Required Package (11)
Service Design and Implementation Anomalies (SD&IA) (174) Code Design Anomaly (125) Poor Code Readability (54)
Messy Code (35)
System Design Anomaly (44) Wrong Dependencies Chain (16)
Lack of ASRs (5)
Poor Security Management (PSM) (126) Coding Level (55) Unsafe Code (23)
Malformed Input (12)
Communication Level (42) Security Dependencies (8)
Wrong Access Control (5)
Application Level (29) Insecure Configuration Management (17)
Missing Security Features (7)
Insufficient Resources (IR) (93) Memory Issue (68) Limited Memory for Process Execution (66)
IDE Problem with Memory (2)
Lack of Human Resources,Tools and Platforms (25) Lack of Tool Support (7)
Deployment Platform Problem (5)
Fragile Code (FC) (24) Poor Implementation of Code (20) Poor Object Oriented Design (6)
Poor Code Reusability (5)
Poor Code Flexibility (11) Tightly Services Components (7)
Divergent Changes (9)
Key Findings of RQ2: We found that not all the issue discussions provide the information about their causes, and finally we identified 2,225 issue discussions containing information about the causes with 31 causes mentioned by the interviewees and 11 causes indicated by the survey participants, which are 2,267 causes instances in total. The cause taxonomy of microservices issues consists of 8 categories, 26 subcategories, and 228 types. The majority of causes are related to General Programming Errors (37.93%), Missing Features and Artifacts (17.02%), and Invalid Configuration and Communication (16.85%).
3.3 Solutions of Issues (RQ3)
The taxonomy of solutions for microservices issues is provided in Table 4. It is worth mentioning that not all the issue discussions provide the information about their solutions. Therefore, we identified 1,899 issue discussions containing the information about the solutions. The taxonomy of solutions is derived by mining developer discussions (i.e., 1,899 solutions), conducting practitioner interviews (i.e., 36 solutions, see Section 2.2.2), and conducting a survey (i.e., one instance of solution, see Section 3.4). Hence, we got a total of 1,936 solutions. We identified a total of 196 types of solutions that can be classified in 8 categories and 35 subcategories. Due to space limitations, we only list the top two types of solutions for each subcategory in Table 4. The detail of the types of solutions can be found in the dataset [19]. The results show that Fix Artifacts (1056 out of 1936), Add Artifacts (360 out of 1936), and Modify Artifacts (210 out of 1936) are the top three categories of solutions. Each solution category is briefly discussed below.
1. Fix Artifacts (1056/1936, 54.54%): During the analysis of developer discussions about microservices issues, we identified that most developers did not explicitly mention any solution to the problems. They fixed the issue in the local repository and send the fixed code (e.g., through a pull request) to the maintainer of the public repository. In this case, from the developer discussions, we could not find exactly what they added, removed, or modified in the project to fix a specific issue. Therefore, we named this category Fix Artifacts. We identified and classified 25 types of fix artifacts solutions in 4 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
  • Fix Code Issue (912, 47.10%): The solutions in this subcategory are related to the direct repair of the source code by developers. We identified 14 types of solutions in this subcategory and among them fix source code of issues, fix illegal symbols (syntax) in code, and clean code are the top three types of solutions.
  • Fix Testing Issue (107, 5.52%): This subcategory covers the types of solutions with which testing issues have been fixed. We collected 2 types of solutions in this subcategory that are debug code and add test services in containers.
  • Fix Build Issue (33, 1.70%): Build systems are essential for developing, deploying, and maintaining microservices systems. In contrast, build failures frequently occur across the development life cycle, bringing non-negligible costs in microservices system development. We collected 3 types of solutions, including fix errors in build files, correct build type, and hide fail status.
  • Fix GUI Issue (4, 0.20%): We collected two types of solutions related to repairing the graphical user interface of microservices systems, and they are fix backwards incompatible ui and fix broken screenshot.
2. Add Artifacts (360/1936, 18.59%): This category covers the types of solutions for addressing missing features, packages, files, variables, and documentation and tool support issues. The interview participants also mentioned that they adopted solutions for addressing several types (e.g., CI/CD, Security) of issues that have occurred because of missing features, and one representative quotation is depicted below.
[…]
To address these issues, we recently adopted the ‘service mesh’ approach. This approach helps address security, latency and scalability, fault identification, and runtime error detection in microservices systems. A service mesh pattern can also provide features for a service health check with the lowest latency to address errors and fault identification issues(P7, Software Engineer).
We identified and classified 33 types of solutions in 10 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
  • Add Features and Services (131, 6.76%): Developers added required features and services to microservices systems. System features and services refer to a process that accepts one or more inputs and returns outputs for particular system functionality [35]. We identified 5 types of solutions in this subcategory, and among them, add missing features, add security features, and add communication protocols are the top three types of solutions.
  • Add Files, Templates, and Interfaces (39, 2.01%): This subcategory covers the solutions for adding missing files, templates, and interfaces in microservices systems. We identified 3 types of solutions, including add files, add templates, and add interfaces.
  • Add Methods and Modules (33, 1.70%): We identified the solutions in which developers mainly added or repaired the missing methods and modules to address several types (e.g., Compilation, Service Execution and Communication, Build) of issues in microservices systems. We identified 3 types of solutions, including add constructor, add parameters in methods, and add security certificates.
  • Add Test Cases (28, 1.44%): We identified the solutions in which developers added test cases to address testing issues of microservices systems. We identified 4 types of solutions in this subcategory in which the top three are add test cases to validate service, generate correct test cases, and add test cases to validate data type.
  • Add Classes and Packages (27, 1.39%): This subcategory covers the solutions in which developers added classes and packages to address the microservices issues. We identified 3 types of solutions, including add packages, add properties, and add objects.
  • Implement Patterns and Strategies (23, 1.18%): The solutions in this subcategory are based on interviewees’ feedback in which they mentioned 19 MSA patterns and strategies for addressing microservices design issues. The most frequently mentioned patterns and strategies are service mesh architecture, service instance per container, and serverless deployment.
  • Add Data Types, Identifiers, and Loops (21, 1.08%): This subcategory contains the solutions for addressing errors at the program initialization level. We identified 5 types of solutions in which the top three are add identifiers, add data types, and add query parameters.
  • Add Dependencies and Metrics (7, 0.36%): The issues related to missing metrics (e.g., monitoring) and dependencies can be resolved by adding the required metrics and dependencies in the microservices systems. We identified 3 types of solutions, including add monitoring metrics, add dependencies, and add stack trace.
  • Add APIs, Namespaces, and Plugins (7, 0.36%): APIs, namespaces, and plugins are the essential part of any microservices systems. This subcategory covers the solutions for addressing missing APIs, namespaces, and plugins. We identified three types of solutions in this subcategory, including add apis, add namespaces, and add plugins.
  • Add Logs (4, 0.20%): Logging activity is specifically related to monitoring microservices systems. We identified 4 types of solutions that have been used to address monitoring issues (e.g., missing logs), in which the top three are add trace id, add trace logging, and add header logging.
3. Modify Artifacts (212/1936, 10.95%): Besides adding new artifacts to the existing system to address the microservices issues, we also identified a large number of solutions in which the developers explicitly mentioned how they modified modules, services, packages, APIs, scripts, methods, objects, data types, identifiers, databases, and documentation to address the microservices issues. We identified and classified 31 types of solutions in 5 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
  • Modify Methods and Objects (78, 4.03%): We found that most developers corrected the proprieties (e.g., method calls, operations, parameters) associated with methods and objects of classes to address various types of issues (e.g., code smells, excessive literals) in microservices systems. We identified 13 types of solutions in this subcategory, among them correct method definition, redefine method operations, and correct method parameters are the top three types of solutions.
  • Modify Packages, Modules, and Documentation (64, 3.30%): This subcategory covers the packages, modules, and documentation that are modified to address the microservices issues. We identified 4 types of solutions in this subcategory, and among them improve documentation (“Resolve ’Create Stack’ issue and update docs, #336”) and update packages are the two leading types of solutions.
  • Modify Data Types and Identifiers (51, 2.63%): This subcategory covers the data types and identifiers that are modified to address the microservices issues. We collected 5 types of solutions, and among them correct naming, correct data types, and correct nil value are the top three types of solutions.
  • Modify APIs, Services, and Scripts (13, 0.67%): This subcategory covers the solutions in which APIs, services, and scripts are modified to address the microservices issues. We identified 6 types of solutions in this subcategory, and among them update scripts, update syntax, and modify services are the top three types of solutions.
  • Modify Databases (6, 0.31%): This subcategory covers the solutions in which database query strings and tables are modified to address the microservices issues. We identified 2 types of solutions related to this subcategory, including modify query strings and modify database tables.
4. Remove Artifacts (39/1936, 2.01%): This category covers the solutions in which artifacts are removed to address several types of microservices issues. We identified and classified 14 types of solutions in 5 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
  • Remove Data Types, Methods, Objects, and Plugins (20, 1.03%): This subcategory covers the solutions in which data types, methods, objects, and plugins are removed to address the microservices issues. We identified 4 types of solutions in this subcategory, including remove data types, remove methods, remove conflicting plugins, and remove objects.
  • Remove Dependencies and Databases (8, 0.41%): This subcategory covers the solutions in which dependencies and database images are removed to address the microservices issues. We identified 2 types of solutions in this subcategory, including remove conflicting dependencies and remove database images.
  • Remove Logs (5, 0.25%): Logging is required to track the communication and identify the failure in microservices systems. This subcategory covers the solutions in which log messages and transaction IDs are removed to address the microservices issues. We identified 3 types of solutions in this subcategory to address the microservices issues, including eliminate log messages, remove transaction id for logging, and unregister from registry.
  • Remove Documentation (5, 0.25%): Several microservices issues were addressed by removing unnecessary or wrong information from project documentation. We identified 2 types of solutions in this subcategory, including remove unnecessary information and remove empty tags.
5. Manage Infrastructure (160/1936, 8.26%): This category captures the solutions that are based on efficient resource utilization for addressing the microservices issues. We identified and classified 20 types of solutions in 2 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
  • Manage Storage (136, 7.02%): Typical microservices systems store their data in dedicated databases for each service. We found several microservices issues that occurred due to the lack of data storage. We identified 9 types of solutions in this subcategory, in which the top three are allocate storage , clean cache, and extend memory.
  • Manage Networking (24, 1.23%): Networking is complicated in microservices systems due to managing an explosion of service connections over the distributed network. We collected 10 types of solutions in which the top three are change proxy settings, server resource management, and disable server groups.
6. Manage Configuration and Execution (50/1936, 2.58%): Managing configuration and execution enables developers to track changes in microservices systems and their consuming applications over time, for example, the ability to track the version history of configuration changes for multiple instances of microservices systems. We identified and classified 13 types of solutions in 2 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
  • Manage Execution (29, 1.49%): This subcategory collects the solutions for issues related to managing commands for executing and configuring microservices systems. We identified 5 types of solutions in this subcategory, and among them execution and configuration management and execute multiple commands are the top two types of solutions.
  • Manage Configuration (21, 1.08%): Managing configurations of each microservice and their instances separately is a tedious and time-consuming task. In this subcategory, we identified 8 types of solutions, and the top three types of solutions are documentation for configuration management, change configuration files, and correct uuid.
7. Upgrade Tools and Platforms (47/1936, 2.42%): The updates of used tools and platforms to develop and manage microservices systems help developers address several issues due to old or legacy versions and protect the microservices systems from security breaches. The interviewees also mentioned a few solutions to address the microservices issues with the help of tools, and one representative quotation is mentioned below.
[…]
Normally, we adopt the solutions according to the type of issues. It could include adding several tools, importing different packages, and adopting successful practices. For example, we improve security by using several open-source API gateways such as OKTA, Spring Cloud gateway, JWT token, and Spring Security. To address service communication issues, we mainly use different ways of communication according to the needs of projects, such as Kafka, RabbitMQ, and Service Mesh. In addition, we automated our continuous integration and delivery process using AWS Code Pipeline. AWS Code Pipeline automates the project release process’s build, test, and deployment phases(P1, Software Architect, Developer).
We identified and classified 21 types of solutions in two subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
  • Upgrade Deployment, Scaling, and Management Platforms (39, 2.01%): This subcategory gathers the solutions for addressing the microservices issues related to the deployment, scaling, and management of CI/CD tools and platforms. We identified 16 types of solutions, and the top 3 types of solutions are upgrade container logging, upgrade load balancer, and upgrade docker files.
  • Upgrade Development and Monitoring Tool Support (8, 0.41%): Development and monitoring are crucial tasks for developers due to the distributed nature of microservices systems. We identified 5 types of solutions in this subcategory, and the top three types of solutions are upgrade kafka flags, upgrade zipkin thrift, and disable tracking.
8. Import/Export Artifacts (12/1936, 0.61%): We found several issues that can be fixed by importing and exporting various artifacts. We identified and classified 3 types of solutions in 2 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
  • Import Artifacts (9, 0.46%): This subcategory gathers the solutions related to importing packages and libraries. We identified 2 types of solutions in this subcategory, which are import packages, and import libraries to address the scalability issues in microservices systems.
  • Export Artifacts (3, 0.15%): In this subcategory, we identified only one type of solutions export packages to fix the Configuration issues in microservices systems (e.g., conflict on configuration file names).
Table 4: Taxonomy of solutions of issues in microservices systems
Category of Solutions Subcategory of Solutions Types of Solutions
Fix Artifacts (1056) Fix Code Issue (912) Fix Source Code of Issues (791)
Fix Illegal Symbols (Syntax) in Code (79)
Fix Testing Issue (107) Debug Code (105)
Add Test Services in Containers (2)
Fix Build Issue (33) Fix Errors in Build File (17)
Correct Build Type (8)
Fix GUI Issue (4) Fix Backwards Incompatible UI (3)
Fix Broken Screenshot (1)
Add Artifacts (360) Add Features and Services (144) Add Missing Feature (114)
Add Security Feature (26)
Add Interfaces, Templates, and Files (39) Add File (37)
Add Templates (1)
Add Methods and Modules (33) Add constructor (24)
Add Parameters in Method (8)
Add Security Certificates (1)
Add Classes and Packages (27) Add Packages (15)
Add Properties(11)
Add Test Cases (28) Add Test Cases to Validate Service (21)
Generate Correct Test Case (3)
Implement Patterns and Strategies (23) Service Mesh Architecture (2)
Serverless Deployment (2)
Add Data Types, Identifiers, and Loops (21) Add Identifiers (11)
Add Data Types (5)
Add Dependencies and Metrics (7) Add Monitoring Metrics (3)
Add Dependencies (2)
Add APIs, Namespaces, and Plugins (7) Add APIs (4)
Add Namespaces (2)
Add Logs (4) Add Trace ID (1)
Add General Purpose Logger (1)
Modify Artifacts (212) Modify Methods and Objects (78) Correct Method Definition (51)
Redefine Method Operations (8)
Modify Package, Module, and Documentation (64) Improve Documentation (56)
Update Packages (6)
Modify Data Types and Identifiers (51) Correct Naming (29)
Correct Data Types (8)
Modify APIs, Services, and Scripts(13) Update Scripts (5)
Update Syntax (3)
Modify Database (4) Modify Database Tables (2)
Modify Query Strings (3)
Remove Artifacts (38) Remove Data Types, Methods, Objects, and Plugins (20) Remove Conflicting Dependencies (17)
Remove Database Images (1)
Remove Dependencies and Databases (8) Remove Conflicting Dependencies (7)
Remove Code Dependencies (1)
Remove Logs (5) Eliminate Log Messages(3)
Remove Transaction ID for Logging(1)
Remove Documentation (5) Remove Unnecessary Information (3)
Remove Empty Tags (1)
Manage Infrastructure (160) Manage Storage (82) Allocate Storage (77)
Clean Cache (5)
Manage Networking (24) Change Proxy Setting (8)
Server Resource Management (5)
Manage Configuration and Execution (50) Manage Execution (29) Manage Configuration Commands (24)
Execute Multiple Commands (2)
Manage Configuration (21) Documentation for Configuration Management (7)
Change Configuration Files (4)
Upgrade Tools and Platforms (43) Upgrade Deployment, Scaling, and Management Platforms (35) Upgrade Container Logging (6)
Upgrade Load Balancer (4)
Upgrade Development and Monitoring Tool Support (8) Upgrade Kafka Flags (2)
Upgrade Zipkin Thrift (2)
Import/Export Artifacts (12) Import Artifacts (9) Import Packages (5)
Import Libraries (4)
Export Artifacts (3) Export Packages (3)
Key Findings of RQ3: We found that not all the issue discussions provide the information about their solutions, and finally we identified 1,899 issue discussions containing information about the solutions with 36 solutions mentioned by the interviewees and 1 solution mentioned by the survey participants, which are 1,936 solutions in total. The solution taxonomy consists of 8 categories, 32 subcategories, and 177 types of solutions. The majority of solutions are related to Fix Artifacts (54.54%), Add Artifacts (18.59%), and Modify Artifacts (10.95%).
3.4 Practitioners’ Perspective (RQ4)
[…]
Practitioners’ Perspective on Issues: We asked the microservices participants which issue they faced while developing microservices systems. The majority of the respondents mentioned that they face all of the issues while developing microservices systems. The practitioners’ responses presented in Table 5 show that IC1: Technical Debt (46.67% Very Often, 24.00% Often), IC2: Continuous Integration and Delivery Issue (26.67% Very Often, 42.67% Often), and IC5: Security Issue (18.67% Very Often, 64.00% Often) occur most frequently than other categories of issues. Some practitioners also suggested 6 types of issues (with 9 instances of issues) that were not part of the initial taxonomy of issues in microservices systems. We added microservices practitioners’ suggested types of issues in Figure 4 and the Issue Taxonomy sheet in [19]. One representative quotation about other types of issues is depicted below.
[…]
In my understanding, there should be some other issues such as lack of (i) understanding of the implementation domain, (ii) defined process for designing, developing, and deploying the projects, and (iii) expertise in programming languages for developing microservices systems(Application Developer, SP2).
Practitioners’ Perspective on Causes: Our survey participants also confirmed the identified causes that could lead to issues in microservices systems. The practitioners’ responses presented in Table 5 show that CC1: General Programming Error (41.33% Strongly Agree, 38.67% Agree), CC8: Fragile Code (23.33% Strongly Agree, 46.67% Agree), and CC2: Missing Features and Artifacts (12.67% Strongly Agree, 55.33% Agree) are the top three cause categories. Some practitioners also suggested 5 types of causes (with 11 instances of causes) that were not part of the initial taxonomy of causes in microservices systems. We added microservices practitioners’ suggested types of causes in Table 5 and the Cause Taxonomy sheet in [19]. One representative quotation about suggested types of causes is depicted below.
[…]
Several microservices issues can be occurred because of (i) separate physical database, (ii) lack of resilience support for the whole application, (iii) excessive tooling, (iv) lack of CI/CD (e.g., DevOps) culture in organizations, and (v) lack of practitioners in teams who have multiple skills(Business Analyst, SP120)
Practitioners’ Perspective on Solutions: We also recorded the practitioners’ perspective about the solutions for the issues occurring during microservices system development. The practitioners’ responses presented in Table 5 show that SC1: Add Artifacts (42.00% Strongly Agree, 39.33% Agree), SC3: Modify Artifact (27.33% Strongly Agree, 37.33% Agree), and SC6: Manage Configuration and Execution (18.00% Strongly Agree, 48.67% Agree) and are the top three solution categories that have been used to address the microservices issues. One practitioner also suggested one type of solution (with one instance of solution) that was not part of the initial taxonomy of solutions in microservices systems. We added microservices practitioners’ suggested types of solutions in Table 5 and the Solution Taxonomy sheet in [19]. One representative quotation about the suggested solutions is depicted below.
[…]
Regular training of the employees on the latest technologies and cloud platforms for developing and managing microservices systems can address several types of security, communication, and deployment issues” (DevOps & Cloud Engineer, SP92).
Table 5: Practitioners’ perspective (in %) on the issue, cause, and solution categories in microservices systems
Microservices Issues (VO-Very Often, O-Often, S-Sometimes, R-Rarely, N-Never)
Categories VO O S R N Mean
Technical Debt 46.67 24.00 12.67 5.33 1.33 3.93
Continuous Integration and Delivery (CI/CD) Issue 26.67 42.67 15.33 7.33 8.00 3.73
Exception Handling Issue 14.67 29.33 27.33 13.33 14.67 3.14
Service Execution and Communication Issue 15.33 44.67 22.67 10.00 7.33 3.51
Security Issue 18.67 64.00 8.67 6.00 2.67 3.90
Build Issue 15.33 43.33 23.33 10.00 8.00 3.48
Configuration Issue 14.00 38.67 25.33 10.67 11.33 3.33
Monitoring Issue 12.00 40.67 22.67 16.67 6.67 3.31
Compilation Issue 16.00 34.67 15.33 14.00 20.00 3.13
Testing Issue 26.67 32.67 20.00 8.00 11.33 3.51
Documentation Issue 13.33 42.00 14.67 11.33 18.67 3.20
Graphical User Interface (GUI) Issue 7.33 30.67 22.00 16.67 20.67 2.79
Update and Installation Issue 10.67 44.00 20.67 12.00 12.00 3.27
Database Issue 18.00 40.67 18.00 16.00 7.33 3.46
Storage Issue 12.00 38.67 22.00 17.33 10.00 3.25
Performance Issue 17.33 46.00 14.67 14.67 6.67 3.51
Networking Issue 12.00 34.67 22.00 20.00 10.67 3.15
Typecasting Issue 12.67 34.00 21.33 10.67 18.00 3.03
Organizational Issue 18.67 64.00 8.67 6.00 2.67 3.90
Causes of Issues (SA-Strongly Agree, A-Agree, UD-Undecided, D-Disagree, SD-Strongly Disagree)
Categories SA A U D SD Mean
General Programming Error 41.33 38.67 12.00 6.00 2.00 4.11
Missing Features and Artifacts 12.67 55.33 25.33 4.67 2.67 3.70
Invalid Configuration and Communication Problem 18.00 47.33 20.00 10.00 4.67 3.64
Legacy Versions, Compatibility, and Dependency Problem 12.67 41.33 27.33 13.33 5.33 3.43
Service Design and Implementation Anomaly 19.33 48.00 18.67 8.67 5.33 3.67
Poor Security Management 18.67 42.67 21.33 10.00 7.33 3.55
Insufficient Resources 16.00 46.00 22.67 10.00 5.33 3.57
Fragile Code 23.33 46.67 12.67 12.67 4.67 3.71
Solutions for Issues (SA-Strongly Agree, A-Agree, UD-Undecided, D-Disagree, SD-Strongly Disagree)
Categories SA A U D SD Mean
Add Artifacts 42.00 39.33 8.67 6.00 3.33 4.11
Remove Artifacts 16.00 60.67 14.00 6.67 2.00 3.83
Modify Artifact 27.33 37.33 25.33 6.00 3.33 3.80
Manage Infrastructure 14.00 48.00 23.33 10.00 4.00 3.58
Fix Artifacts 18.00 46.67 24.00 6.00 4.67 3.68
Manage Configuration and Execution 18.00 48.67 18.67 9.33 4.67 3.66
Upgrade Tools and Platforms 15.33 52.00 22.00 8.00 2.00 3.71
Import/Export Artifacts 14.67 50.00 28.67 3.33 2.67 3.71
Statistical significance on the issue, cause, and solution categories in microservices systems: We analyzed the practitioners’ responses across one pair of demographic groups in Table 6. The first column of Table 6 lists the categories of issues, causes, and solutions presented to the survey participants. The subsequent columns of Table 6 show the Likert Distribution, Mean, p-value, and Effect Size. Analyzing survey data using the Likert Distribution, Mean, p-value, and Effect Size provides insights into response patterns, central tendency, statistical significance, and practical significance, respectively. The practitioners’ responses are grouped into Experience $\leq$ 6 Years vs. Experience > 6 Years group to check the test whether there are statistical differences between the two groups on the same variable or not.
[…]
Observations: We made the following observations based on the practitioners’ responses.
  • There are no major statistically significant differences in practitioners’ responses on the issue, cause, and solution categories in microservices systems.
  • The observed statistically significant differences between experienced-based grouping indicate that the experience of microservices practitioners does not affect the survey responses.
  • The survey findings indicate that most issues are related to Technical Debt, CI/CD, and Security; most causes are associated with General Programming Errors, Fragile Code, and Missing Features and Artifacts; and most solutions are associated with Add Artifacts, Modify Artifacts, and Manage Configuration and Execution.
  • More than 50% of the respondents indicated that they frequently (Very Often and Often) encountered different issues belonging to the given list of issue categories.
  • Our results indicate that many practitioners rarely or never face GUI (16.67% Rarely, Never 20.67%), Compilation (14.00% Rarely, 20.00% Never), and Networking issues (20.00% Rarely, 10.66% Never) in microservices system development.
Key Findings of RQ4: (i) There are no major statistically significant differences in the issue, cause, and solution categories in practitioners’ responses, (ii) practitioners frequently face Technical Debt, CI/CD, and Security issues, (iii) most causes are associated with General Programming Errors, Fragile Code, and Missing Features and Artifacts, and (iv) most solutions are associated with Add Artifacts, Modify Artifact, and Manage Configuration and Execution. The survey participants generally confirmed the categories of issues, causes, and solutions derived from mining developer discussions in open-source microservices systems. However, the survey participants also indicated several other types of issues, causes, and solutions that help improve our taxonomies.
Table 6: Statistical significance on the issue, cause, and solution categories in microservices systems
Categories # Likert Distributions Experience Based Grouping
In Total Mean P-value Effect Size
Microservices Issues
Technical Debt IC1 [Uncaptioned image] 3.89 0.47 0.45
Continuous Integration and Delivery (CI/CD) Issue IC2 [Uncaptioned image] 3.73 0.35 0.30
[…]
Service Execution and Communication Issue IC4 [Uncaptioned image] 3.51 0.17 0.35
Security Issue IC5 [Uncaptioned image] 3.90 0.67 0.16
Build Issue IC6 [Uncaptioned image] 3.48 0.14 0.12
Configuration Issue IC7 [Uncaptioned image] 3.33 0.14 -0.06
Monitoring Issue IC8 [Uncaptioned image] 3.31 0.14 0.23
[…]
Testing Issue IC10 [Uncaptioned image] 3.51 0.09 -0.36
Documentation Issue IC11 [Uncaptioned image] 3.20 0.02 0.31
Graphical User Interface (GUI) Issue IC12 [Uncaptioned image] 2.79 0.08 0.60
[…]
Database Issue IC14 [Uncaptioned image] 3.46 0.25 0.03
Storage Issue IC15 [Uncaptioned image] 3.25 0.12 0.31
Performance Issue IC16 [Uncaptioned image] 3.51 0.21 -0.02
[…]
Typecasting Issue IC18 [Uncaptioned image] 3.03 0.12 0.38
Organizational Issue IC19 [Uncaptioned image] 3.52 0.14 0.11
Causes of Issues
General Programming Error CC1 [Uncaptioned image] 4.11 0.53 0.17
Missing Features and Artifacts CC2 [Uncaptioned image] 3.70 0.47 -0.04
Invalid Configuration and Communication Problem CC3 [Uncaptioned image] 3.64 0.25 -0.31
Legacy Versions, Compatibility, and Dependency Problem CC4 [Uncaptioned image] 3.43 0.25 0.20
Service Design and Implementation Anomaly CC5 [Uncaptioned image] 3.67 0.35 0.16
Poor Security Management CC6 [Uncaptioned image] 3.55 0.21 -0.20
Insufficient Resources CC7 [Uncaptioned image] 3.57 0.40 0.02
Fragile Code CC8 [Uncaptioned image] 3.71 0.25 -0.12
Solutions for Issues
Add Artifacts SC1 [Uncaptioned image] 4.11 0.40 0.20
Remove Artifacts SC2 [Uncaptioned image] 3.83 0.60 -0.01
Modify Artifact SC3 [Uncaptioned image] 3.80 0.21 0.10
Manage Infrastructure SC4 [Uncaptioned image] 3.58 0.30 -0.12
Fix Artifacts SC5 [Uncaptioned image] 3.68 0.30 0.20
Manage Configuration and Execution SC6 [Uncaptioned image] 3.66 0.35 0.12
Upgrade Tools and Platforms SC7 [Uncaptioned image] 3.71 0.35 -0.27
Import/Export Artifacts SC8 [Uncaptioned image] 3.71 0.47 0.22
4 Discussion
We analyze in this section the relationship between the issues, causes, and solutions (Section 4.1) followed by presenting the implications of the research (Section 4.2).
4.1 Relationship between issues, causes, solutions
While the taxonomy in Figure 4 provides a categorization of issues in microservices systems, a mapping between the issues, their causes, and solutions is presented in Figure 5. Mapping diagrams are frequently used in systematic mapping studies—relying on bubble plots—to correlate data or concepts along different dimensions [36]. We have chosen the mapping diagram to depict the issues on the Y-axis and their causes on the X-axis, and to present the solutions that can address these issues at the intersections of the X and Y axes. Eight different types of symbols, labeled at the top of the figure, are used to represent various types of solutions. The interpretation of the mapping in Figure 5 involves locating a given issue on the Y-axis, associating this issue with its cause(s) on the X-axis, and identifying the possible solutions to address the issue, as elaborated and exemplified below. Figure 5 indicate that it is possible that one issue could cause other issues. Similarly, one cause may be the reason for several issues, or several causes may contribute to one issue.
  • Issues in microservices systems (Y-axis): A total 19 categories of issues, adopted from Figure 4, are presented on the Y-axis. For example, one of the issue categories Technical Debt has a total of 687 instances of issues as presented in Figure 4.
  • Causes of the issues (X-axis): A total of 8 categories of causes are presented on the X-axis, mapped with the corresponding issues. For example, General Programming Errors, such as incorrect naming and data type (157, 6.92%), testing error (25, 1.10%), and poor documentation (22, 0.97%), are the predominant causes of Technical Debt issues. In comparison, the causes like Poor Security Management have no impact on Technical Debt issues.
  • Solutions to resolve the issues: A total of 8 categories of solutions are presented at the intersection of the issues and their causes. For example, solutions, such as fixing an artifact (code, GUI, errors in build files) or removing an artifact (code dependency, empty tag, unnecessary documentation) can help with fixing a majority of the Technical Debt issues.
The mapping in Figure 5 can have a diverse interpretation based on the intent of the analysis that may include but is not limited to frequency analysis and data correlations. It is virtually impossible to elaborate on all possible interpretations, however; to exemplify some of the possible interpretations can be as follows:
  • What are the most and least frequently occurring issues in microservices systems? As per the mapping, the most frequent (top 3) issues are related to Technical Debt, CI/CD, and Exception Handling, representing a total of 764 identified issues. On the other hand, the least frequent issues relate to Organizational, Update and Installation, and Typecasting issues.
  • What are the most and least common causes of a specific category of issues? The mapping of the causes suggests that General Programming Errors, Invalid Configuration and Communication Problems along with Legacy Versions, Compatibility and Dependency Problem are the most common causes of the issues. Similarly, Fragile Code represents the least common cause for issues in microservices systems.
  • What are the most and least recurring solutions to fix a specific category of issues? Fixing Artifacts, Add Artifacts, and Modify Artifacts represent the most recurring solutions to address microservices issues, whereas importing/Export Artifacts, Upgrade Tools and Platforms, and Manage Configuration and Execution categories represent the least recurring solutions to address microservices issues.
Figure 5: Mapping between issues, causes, and solutions in microservices systems
4.2 Implications
[…]
Practitioners should prioritize managing technical debt, including code debt and service design debt, in microservices systems to mitigate potential risks and ensure the long-term health of the systems. This involves addressing source code issues like refactoring and duplication, and adhering to successful practices in microservices system design.
[…]
Practitioners should pay attention to correct configuration settings, updating and managing dependencies, and keeping abreast of best practices in tools and platforms being used (like Docker, Kubernetes, AWS, Google Cloud). They should also consider investing in upgrading tools and platforms, and fixing artifacts as a common resolution to many CI/CD issues. This will ensure a smoother and more reliable CI/CD process.
[…]
To ensure robust security in microservices systems, practitioners should implement multi-layered security solutions. This involves strategies for data centers, virtualization, communication, and orchestration. Measures such as secure certificates, encryption, and access control policies are essential for achieving strong security management in microservices.
[…]
The majority of build errors are related to build scripts, plugin compatibility, and docker build failures. It is crucial for practitioners to focus on these aspects and ensure that build scripts are error-free, plugins are compatible, and Docker build files do not contain unnecessary dependencies.
[…]
Practitioners need to be cautious of how microservices communicate with each other. Poor implementation of communication between microservices can lead to insecure communication, latency, lack of scalability, and difficulties in fault identification at runtime. By focusing on effective communication protocols and configurations, practitioners can ensure smoother service interactions and mitigate the challenges related to security, performance, and fault tolerance.
[…]
In microservices systems, proper configuration is vital due to the extensive service instances and third-party integrations. Practitioners must carefully manage configurations to avoid increased latency and slow microservice calls. This involves diligently handling settings and files, and avoiding errors such as configuration mismatches and incorrect file paths, especially during development, implementation, and deployment phases.
[…]
The dynamic and distributed nature of microservices systems necessitates robust monitoring infrastructures. Since traditional tools may not suffice due to their focus on specific components or overall health, it is crucial for practitioners to employ a combination of tools and strategies that effectively monitor various aspects such as errors, faults, performance issues, and the health of the system across distributed environments.
[…]
Testing in microservices systems is complex and multifaceted, and practitioners must be cognizant of specific challenges, such as coordinating tests among multiple services, ensuring adequate test coverage, and dealing with polyglot code bases and network communications. Special attention should be paid to issues like faulty test cases, debugging, and load test cases.
[…]
Practitioners should explore dynamic storage platform assignment techniques to optimize storage utilization. By assigning storage platforms (e.g., containers or virtual machines) based on microservice requirements, practitioners can improve resource allocation, enhance system performance, and efficiently manage energy, cost, and compliance considerations.
[…]
The distributed nature of microservices systems leads to distributed ownership of databases, with each microservice having its private data store. This distribution poses challenges related to database implementation, data accessibility, and database connectivity. Practitioners need to be aware of these challenges and consider strategies to effectively manage databases in microservices architectures.
[…]
Addressing networking issues in microservices systems involves managing network infrastructure and modifying artifacts. Prioritizing infrastructure management, including protocols, ports, topologies, and service accessibility, enables practitioners to mitigate networking issues and ensure smooth operations.
[…]
The decentralized nature of microservices, with numerous independent services deployed on different infrastructures and platforms, can lead to performance overhead and increased resource consumption. Practitioners need to be aware of these challenges and take them into account when designing, deploying, and managing microservices systems. They should consider strategies and techniques to optimize resource utilization and minimize performance overhead to ensure efficient and scalable microservices performance.
[…]
Practitioners in microservices systems should be aware of typecasting issues when using multiple programming languages and technologies. Understanding language differences enables effective anticipation and mitigation of typecasting problems.
5 Threats to Validity
This section reports the potential threats to the validity of this research and its results, along with mitigation strategies that could help minimize the impacts of the outlined threats based on [17]. The threats are broadly classified across internal, construct, external, and conclusion validity.
5.1 Internal Validity
Internal validity examines the extent to which the study design, conduct, and analysis answer to the research questions without bias [17]. We discuss the following threats to internal validity.
Improper project selection. The first internal validity threat to our study is an improper selection of OSS projects for executing our research plan. Mitigation. We used a multi-step project selection approach to control the possible threat associated with subject system selection (see Phase 1 in Section 2.2 and Figure 2). A step-wise and criteria-driven approach for the project selection has been used that helped us to include the relevant (see Table 1) and eliminate irrelevant open-source microservices projects.
Instrument understandability. The participants of interviews and surveys may have a different understanding of the interview and survey instruments. Mitigation. We adopted Kitchenham and Pfleeger’s guidelines for conducting surveys [29]. We also piloted both interview and survey instruments to ensure understandability (see Section 2.3.1 and Section 2.4.1). Our questionnaire for the survey were in English. However, during the interviews, we found that a few participants could not conveniently convey their answers in English. Therefore, we requested them to answer in their native languages, and for the latter we translated the answers into English.
Participants selection and background: The survey participants may not have sufficient expertise to answer questions. Mitigation. We searched for microservices practitioners through personal contacts and relevant platforms (e.g., LinkedIn and GitHub). We explicitly made participants’ characteristics (e.g., roles and responsibilities) in the interview and survey preamble, and we selected only those practitioners with sufficient experience designing, developing, and operationalizing microservices systems. For example, the average experience of the interview participants is 5.33 years, and they mainly have responsibilities for designing and developing microservices systems (see Table 2 and Figure 3).
Interpersonal bias in extracting and synthesizing data. (i) The interpersonal bias in the mining developer discussions and analysis process may threaten the internal validity of the study findings. Mitigation. To address this threat, we defined explicit data collection criteria (e.g., exclude the issues of general questions). We had regular meetings with all the authors in data labelling, coding, and mapping. The conclusions were made based on the final consensus made by all the authors. (ii) The survey and interview participants may be slightly biased in providing actual answers due to company policies, work anxiety, or other reasons. Mitigation. To mitigate this threat, we highlighted the anonymity of the participants and their companies in the interview and survey instruments preamble. (iii) The interviewer of the study may be biased toward getting the favourite answers. Mitigation. We sent our interview questions 3 to 4 days before the interview to the interviewees. Hence, they had sufficient time to understand the context of the study and could provide the required feedback on the completeness and correctness of the developed taxonomies.
5.2 Construct Validity
Construct validity focuses on whether the study constructs (e.g., interview protocol, survey questionnaire) are correctly defended [17]. Microservices issues, causes, and solutions are the core constructs of this study. Having said this, we identified the following threats.
Inadequate explanation of the constructs: This threat refers to the fact that the study constructs are not sufficiently described. Mitigation. To deal with this threat, we prepared the protocols for mining developer discussions and conducting interviews and surveys, and these protocols were continuously improved during the internal meetings, feedback, and taxonomy refinements. Mainly, the authors had meetings (i) to establish a common understanding of issues, causes, and solutions (see Figure 1), (ii) for extracting the required data to answer the RQs, and (iii) for evaluating interview and survey question format, understandability, and consistency. We also invited two survey-based research experts to check the validity and integrity of the survey questions. Based on their feedback, we included Figure 4, Table 3, and Table 4 as a part of the interview and survey questions for improving the participants’ understandability of the issues, their causes and solutions in microservices systems.
Data extraction and survey dissemination platforms: This threat refers to the authenticity and reliability of the platforms we used for data collection and survey dissemination. Mitigation. (i) To mitigate this threat for data collection, we identified 2,641 issues from the issue tracking systems of 15 open-source microservices systems on GitHub which were confirmed by the developers. (ii) We disseminated the survey at social media and professional networking groups. The main threat to survey dissemination platforms is identifying the relevant groups. We addressed this threat by reading the group discussion about microservices system development and operations. After ensuring that the group members frequently discussed various microservices aspects, we posted our survey invitation.
Inclusion of valid issues and responses: The threat is mainly related to the inclusion and exclusion of issues from the issue tracking systems and responses from survey participants. Mitigation. We defined explicit criteria for including and excluding issues from the issue tracking systems (see Section 2.2.2) and responses from survey participants (see Section 2.4.2). For example, when screening issues from the issue tracking systems, we excluded those issue discussions consisting of general questions, ideas, and proposals. Similarly, we excluded those responses that were either randomly filled or filled by research students and professors who were not practitioners.
5.3 External Validity
External validity refers to the extent to which the study findings could be broadly generalized in other contexts [17]. The sample size and sampling techniques might not provide a strong foundation to generalize the study results. It is the case for all three data collection methods (mining developer discussions from open-source microservices projects, interviews, and surveys) used in this study. Mitigation. (i) To minimize this threat, we derived the taxonomies of issues, causes, and solutions from a relatively large number of issues from 15 sampled open-source microservices projects belonging to different domains by involving multiple researchers. (ii) The taxonomies have been evaluated and improved by taking the feedback of experienced microservices practitioners through the interviews. (iii) A cross-sectional survey was conducted based on the derived and evaluated taxonomies. Overall, we received 150 valid responses from 42 countries of 6 continents (see Figure 3(a)) having varying experience (from less than one year to more than ten years, see Figure 3(b)) with different roles (see Figure 3(c)), and working with diverse domains (see Figure 3(d)) and programming languages and technologies (see Figure 3(e)) to develop microservices systems. Our study is primarily based on analyzing qualitative data, which presents potential threats, such as subjectivity bias, low reproducibility, data overload, and a lack of transparency that researchers may encounter. Mitigation. To mitigate these threats during data analysis, we employed thematic analysis by following the guidelines (see Section 2.2.2) and involved all the researchers to reduce subjectivity bias, enhance replicability, manage data overload, and ensure transparency.
textitInterview data saturation: This threat arises from the possibility that the 15 interviews we conducted to validate and expand our proposed taxonomies may not be sufficient to achieve data saturation in our study. Mitigation. To address this threat, we conducted interviews with practitioners who had varied backgrounds (i.e., responsibilities, skills in programming languages, work domains, experience, countries) in microservices systems (see Table 2). Second, we asked open-ended questions, which allowed the interviewees to spontaneously express their views on the issues, causes, and solutions (see the Interview Questionnaire sheet in [19]). Third, we presented the interviewees with pre-existing taxonomies that were derived from analyzing developer discussions in 15 open-source microservices systems. We asked the interviewees to pinpoint any issue, cause, and solution they felt absent. Fourth, while we were analyzing interview data, we started observing recurring themes and patterns in the responses, which was an indication of data saturation. As we did not encounter any new significant issues, causes, or solutions after the 15th interview, we decided to stop conducting further interviews. However, we encourage further research involving more interviews or additional data sources to provide more insights and refinement to our findings.
We acknowledge that the findings of this study may not be generalizable, fully replicable, or representative of the issues, causes, and solutions for all types of microservices systems. However, the size of the investigated issues, the number of microservices systems, interviews with microservices practitioners, the survey population, and systematic data analysis methods (e.g., thematic analysis) can partially strengthen the overall generalizability of the study results.
5.4 Conclusion Validity
Conclusion validity is related to dealing with threats that affect the correct conclusions in empirical studies [17]. Concluded findings of the results may be based on a single author’s understanding and experience, and conflicts on the conclusions between authors may not be sufficiently discussed or resolved. Mitigation. To address this threat, the first author extracted and analyzed study data (i.e., from the developer discussions, interviews, and survey). All other authors comprehensively reviewed the data through multiple meetings. Conflicts on data analysis results were resolved through mutual discussions and brainstorming among all the authors. Different researchers can interpret the inclusion and exclusion criteria differently which influences the conclusions of the study. Mitigation. To minimize the effect of this issue, we applied the explicitly defined inclusion and exclusion criteria during the study data screening (see Section 2.2.2 - Step B). Finally, the interpreted results and conclusions have been confirmed by arranging several brainstorming sessions among all the authors.
6 Related Work
We performed a thorough literature search via eight major databases, including Google Scholar, IEEE Xplore, ACM Digital Library, ScienceDirect, Scopus, Web of Science, SpringerLink, and Wiley Online Library, to ensure that no relevant research study has been overlooked during the literature review. The literature search revealed that no study (except our previous work [18]) has empirically investigated the issues faced, causes reported, and solutions adopted by practitioners to develop microservices systems. We discuss the existing research that can be broadly classified into two categories including (i) mining OSS repositories to extract reusable knowledge for MSA (Section 6.1) and (ii) empirical studies on issues in microservices systems and software systems (Section 6.2). A conclusive summary and comparative analysis (Section 6.3) position the proposed research in the context of mining issues from OSS repositories and justify the scope and contributions of this research.
6.1 Mining OSS Repositories to Extract Reusable Knowledge for MSA
6.1.1 Mining (Anti-)Patterns and Tactics to Architect MSA
From an MSA perspective, architectural patterns [34] and tactics [66] represent empirically derived knowledge, best practices, and recurring solutions to address frequently occurring issues during architecture-centric design and development of service-driven systems [10]. To discover patterns and tactics for architecting MSA, a number of studies have focused on analyzing open source repositories – mining version controls, searching change logs, and exploring design documents etc. – to investigate historical data that represents recurring solutions as patterns [10, 67, 68, 69].
The investigation of historical data involves postmortem analysis of development activities (e.g., software refactoring, testing, evolution) [70] as well as the knowledge of architects (e.g., developer discussions, code documentation) [71], which is captured via open source repositories, such as GitHub, Stack Overflow, or customized databases [10, 67, 71]. Specifically, Marquez et al. [10] explored source code artifacts of microservices projects, i.e., configuration files, and framework dependencies, available on GitHub to mine 17 architectural patterns addressing a set of quality attributes. Compared to the design and implementation-specific knowledge for microservices reported in [10, 67], Armin et al. [68] explored the evolution phase from traditional architectures to MSA and identified 15 migration patterns that support the evolution of legacy software towards a modular MSA. Compared to the pattern-based solutions discussed above, architectural tactics represent design decisions that focus on improving one specific quality aspect of MSA, such as service availability and fault avoidance for security critical systems [66]. Amoroso et al. [72] curated and manually labeled a dataset of 378 Dockerized open-source microservice projects from an initial pool of 389,559 candidates. The dataset includes project metadata such as contributors, dependencies, and application purposes, supporting future empirical research on microservice practices and evolution. Lercher et al. [73] investigated API evolution in microservice systems through interviews with practitioners, identifying six strategies and six challenges in handling REST API changes. They proposed automated mechanisms for extracting structural and behavioral API changes and notifying consumer teams. Their findings highlight the socio-technical complexities of API evolution in microservices, complementing prior work on migration and evolvability. Lubas et al. [74] mined GitHub to curate two datasets: one of 553 microservice applications and another of 8 representative workload repositories. Their analysis provides insights into programming languages, frameworks, and architectural components used in practice, supporting benchmarking and empirical studies on workload characterization in microservice systems.
While patterns promote best practices to develop MSA, anti-patterns represent a class of patterns that have been perceived as best practices and commonly used but are proven to be ineffective and/or counterproductive, such as bad smells in service design [14]. Taibi et al. [70] conducted interviews of microservices developers, i.e., accumulating practitioners’ experiences to create a taxonomy of 20 anti-patterns including organizational (team oriented and technology/tool oriented) and technical (internal and communication) anti-patterns. The anti-pattern taxonomy aims to help microservices practitioners to avoid counterproductive patterns and design decisions. Tighilt et al. [75] proposed a specification and detection approach for microservice antipatterns. Their study defined formal detection rules, validated them on open-source systems, and demonstrated their utility for supporting maintenance and quality improvement in microservice architectures. This work goes beyond descriptive taxonomies by providing automated mechanisms for identifying recurring bad practices. In comparison to the repository mining approaches that investigate code and architectural-centric artifacts available on GitHub such as [66, 10, 69], Bandeira et al. [71] reflected a human-centric view of MSA design by analyzing developer discussions on Stack Overflow. Their study classified a total of 1,043 microservice tagged posts into three categories namely technical (44.87%), conceptual (30%), and non-related (25.13%) discussions to explore the processes, tools, issues, solutions, etc. that developers find most exciting or challenging while implementing MSA.
6.1.2 Analyzing Issues for Evolving Systems to MSA
In recent years, research and practices on the migration of legacy systems to microservices systems have gained significant attention with empirical evidence derived from industrial practices on the role of MSA in software evolvability [76, 40], service extraction from legacies [77, 78], and identifying the motivations and challenges for legacy migration to MSA [2]. For instance, Bogner et al. [76] reported the results of 17 structured interviews with 14 microservices practitioners from 10 projects to investigate a multitude of issues, such as tool support, patterns, and process to manage TD and enhance software evolvability using microservices. The findings of the interviews recommended a number of techniques, including but not limited to code review and service slicing, that can address a number of microservices issues related to service granularity, composition, coupling, and cohesion. Lenarduzzi et al. [40] explored the issues of TD when legacy software is migrated to microservices. They investigated four years history of a project having 280K lines of code and concluded that although TD spiked initially due to the development of new microservices, however; after a short period of time the TD grew slower in the microservices system than in the (legacy) monolithic system. Carvalho et al. [78] conducted an online survey with 26 specialists, followed by individual interviews with 7 of them to understand the challenges pertaining to the migration of existing systems to microservices architecture. Their study results highlight that extracting the microservices from legacy components and monolithic source code modules represents the most critical challenge during the re-engineering or migration of legacies toward microservices systems. A number of empirical studies, such as [77, 2], engaged industrial practitioners to understand the processes, motivations, and challenges related to legacy system migration in general and service extraction in particular. The results of these studies provide recommendations and guidelines about service maintainability and scalability as the most important motivations for the enterprise-scale adoption of MSA.
6.2 Empirical Studies on Issues in Microservices Systems and Software Systems
6.2.1 Smells & Performance Issues in MSA Systems
Code smells and architectural smells (a.k.a bad smells) are often synonyms as a microservices anti-pattern reflecting symptoms of poorly designed microservices that decrease code understandability and increase effort for maintainability [70, 14]. According to practitioners’ suggestions and industrial case studies, detecting bad smells in microservices systems is critical for large-scale microservice systems [14]. Walker et al. [79] provided tool support for the automatic detection of bad smells in microservices systems, and the tool MSANose can detect up to eleven microservices specific bad smells within microservices applications using bytecode and/or source code analysis throughout the development process or even before its deployment to production. In addition, a number of other prevalent issues in microservices systems, such as faults, bugs, performance issues, and service decomposition problems, are detailed in [15, 16, 80]. More specifically, Wu et al. [16] proposed a solution named Microservice Root Cause Analysis (MicroRCA), which works by inferring the root causes of performance issues by correlating application performance symptoms with corresponding system resource utilization. Their proposed solution MicroRCA can addresses performance related issues in microservices systems by analyzing resource utilization and throughput of the services.
6.2.2 Taxonomies of Issues & Faults in Software Systems
While exploring issues from the microservices system point of view, it is important not to overlook the most recent taxonomies, empirical studies, and proposed solutions that address a multitude of issues, errors, and faults in non-MSA systems such as, deep learning [81] and application build systems [45]. In particular, a taxonomy of the types of faults in deep learning systems [81] and the types of build issues, their symptoms, and fix patterns [45] inspired our work on microservices issues, causes, and solutions. From the MSA perspective, a recently conducted study [15] identified typical faults occurring in microservices systems, practices of service debugging, and challenges faced by developers while addressing these faults. For example, a fault such as “transactional service failure” is due to overloaded requests to a third-party (payment gateway) service, ultimately leading to denial of service issues. Our proposed research draws inspiration from empirically derived taxonomies [81, 45] and goes beyond issue categorization to investigate their causes and proposed solutions as resolution strategies to fix multi-faceted issues related to Security, Testing, and Configuration that impact architecting and implementing microservices systems.
6.3 Conclusive Summary
The studies reported in [79, 15, 16] are grounded in the empirical analysis of microservices systems to identify a multitude of issues, such as faults, bad smells, and performance issues faced by practitioners while designing, developing, and deploying microservices systems. To complement empiricism in microservices research and development, our proposed study mined the social coding platform (i.e., 15 open-source microservices systems on GitHub) and identified issues faced by developers with improvement and validation by microservices practitioners. While there is work on mining reusable knowledge [10, 67] - patterns and best practices – from software repositories to develop microservices systems, there is no research on mining knowledge, overseeing the broader microservices system development life cycle to streamline the plethora of issues, their causes and fixing strategies. Our study primarily focuses on survey-driven validation of the issues, causes, and solutions of microservices systems by practitioners, and complements the body of research comprising some of the recent industrial studies on evolvability [76], migration [40], and debugging [15] of microservices systems.
We provide a comparative analysis of the most relevant existing research and our proposed study in Table 7 based on (i) focus of investigation: specific aspects to be empirically investigated, (ii) research method: specific approaches employed for investigation, (iii) number of issues: specific number of the identified issues, and (iv) method for evaluation: specific method for evaluation. Additionally, the study reference, year of publication, and contribution of the research study are provided as complementary information for a fine-grained presentation of the results. For example, an empirical study reported in [15] published in 2021 relies on an industrial survey (i.e., applied research method) to investigate common faults and debugging practices in microservices systems (i.e., focus of investigation). The study results (i.e., main contribution) highlight typical faults, state-of-practice on debugging them, and indicate the challenges faced by developers for microservices systems.
Table 7: Comparative analysis and summary of existing research vs. this research
Reference Year Focus of Investigation Research Method # of Issues
[15] 2021 Fault analysis, Debugging Industry survey Not explicit mentioned
Contribution: An industrial survey to learn typical faults, current practice of debugging, and the challenges faced by developers for microservices systems.
[16] 2020 Performance issues MSA application monitoring Not explicit mentioned
Contribution: MicroRCA, as a solution to identify root causes of performance issues in MSAs by monitoring application performance and resource utilization.
[76] 2019 Evolvability analysis Structured interviews 16 types of challenges
Contribution: A qualitative study to explore the tools, metrics, patterns, and practitioners’ view on practices and challenges of microservices system evolvability.
[79] 2020 Code smells Static code analysis 11 types code smells
Contribution: MSANose tool to detect microservice-specific code smells using bytecode and/or source code analysis.
[82] 2023 Security points Repository mining
Contribution: An empirical study that investigates 10 GitHub projects and 306 Stack Overflow posts to identify security challenges and practices for MSA systems.
[83] 2022 Microservice problems and solutions Repository mining 14 problems, 47 solutions
Contribution: An empirical study that analyzes 17522 Stack Overflow posts to explore problems and solutions corresponding to microservice development process.
[84] 2022 Architectural smells Interviews, Repository mining 6 types of architectural smells
Contribution: An industrial case study that collects repository data and practitioners’ views on the impacts, causes, and solutions of architectural smells in MSA.
[75] 2023 Antipattern specification and detection Repository mining, Static analysis Catalog of antipatterns
Contribution: This study formalizes the specification of microservice antipatterns and provides automated detection rules validated on open-source systems to support maintenance and quality assurance.
[72] 2024 Dataset of OSS microservice projects Repository mining, Manual curation 378 projects
Contribution: A curated dataset of 378 Dockerized open-source microservice projects.
[73] 2024 API evolution strategies, challenges, and automation Practitioner interviews (17), Proposed automation 6 strategies, 6 challenges
Contribution: An empirical study on strategies and pitfalls in microservice API evolution, with proposals for automated change extraction and notification.
[74] 2025 Dataset of microservice applications and workloads Repository mining, Empirical analysis 553 applications, 8 workloads
Contribution: Curated two GitHub-based datasets of 553 microservice applications and 8 workload repositories, offering insights into languages, frameworks, and components for benchmarking and workload studies.
Our previous study [18] 2022 Issues and Causes in microservices systems Repository mining 138 types of issues, 109 types of causes
Contribution: This study presents a taxonomy of issues, and identifies and maps to their causes in microservices systems.
This study 2023 Issues, Causes, and Solutions in microservices systems Repository mining (2,641 issues), Structured interviews (15 interviews), Practitioners’ survey (150 responses) 402 types of issues, 228 types of causes, 177 types of solutions
Contribution: This study presents empirically validated taxonomies for issues, causes, and solutions in microservices systems, maps the relationships between them, and provides the dataset for further research.
7 Conclusions
This paper empirically investigates the issues, causes, and solutions of microservices systems by employing a mixed-methods approach. Our study collected data from 2,641 issues from the issue tracking systems of 15 open-source microservices systems on GitHub, 15 interviews and an online survey completed by 150 practitioners. The primary contribution of this work is rooted in the taxonomies of issues, causes, and solutions in microservices systems. The taxonomy of issues consists of 19 categories, 54 subcategories, and 402 types of issues. The taxonomy of causes contains 8 categories, 26 subcategories, and 228 types of causes, whereas the taxonomy of solutions includes 8 categories, 32 subcategories, and 177 types of solutions. Overviewing the results of this study, the major issues for microservices systems are Technical Debt, Continuous Integration and Delivery, and Exception Handling issues. The majority of the issues occur due to General Programming Errors, Missing Features and Artifacts, and Invalid Configuration and Communication. These issues are mainly addressed by Fixing, Adding, and Modifying artifacts.
Based on the results and our observations in this study, we proposed the future directions in three aspects:
New Techniques: Proposing (i) techniques for controlling TD through the design of microservices systems, (ii) multi-layered security solutions for fine-grained security management, (iii) architecting and design techniques for microservices systems with a particular focus on highly resilient and low coupled microservices systems, (ii) solutions to trace and isolate communication issues to increase fault tolerance, (iv) techniques for dynamically optimizing configuration settings of microservices, (v) techniques for identifying critical paths for improving performance and removing performance bottlenecks due to configuration issues, (vi) techniques for detecting security breach points during the configuration of microservices systems, (vii) strategies to automatically test APIs, load, and application security for microservices systems, (viii) techniques for dynamically assigning storage platforms according to the requirements of microservices, (ix) techniques for storing and accessing decentralized and shared data without losing the independence of individual microservices, (x) techniques for optimizing resource utilization (e.g., CPU and GPU usage), (xi) a framework for improving scalability to increase microservices performance, and (xii) techniques for converting code of one language to another used in microservices systems.
Empirically-Grounded Studies and Guidelines: Conducting empirical studies to (i) investigate technical debt at the code, design, and communication level of microservices systems, (ii) explore the efforts required to fix the build issues in microservices systems at different phases (e.g., compilation and linking phases), and (iii) compare the types of problems and the needed efforts to fix the build issues in microservices systems with monolithic systems. We also argue for preparing the strategies and guidelines for (i) addressing cd pipeline errors, (ii) establishing issue, cause, and solution knowledge base for microservices systems in the multi-cloud (e.g., AWS, Google Cloud) containerized environment, (iii) organizing polyglot databases for efficient read and write operations by considering performance, and (iv) addressing microservices security vulnerability and related risks at various levels, such as data centers, cloud providers, virtualization, communication, and orchestration.
Tools: Designing and developing intelligent tools (i) to identify, measure, prioritize, monitor, and prevent TD in microservices systems, (ii) for distributed tracing and real-time performance monitoring, and (iii) to fix Networking issues like localhost, ip address, and webhook errors. Moreover, The outcome of our study (e.g., issue, cause, and solution taxonomies) can help propose a systematic process and develop an intelligent recommendation system to (semi-)automatically identify the issues and causes of microservices system development, as well as recommend the solutions to address those issues. The recommendation system can assist microservices practitioners in efficiently and effectively designing, developing, and operationalizing microservices systems.
In conclusion, this paper has presented an overview of various issues that can occur in microservices systems. We have discussed the challenges and pitfalls in the design and implementation of microservices systems and highlighted the patterns and trends in the types of issues that arise. Additionally, we have provided insights into the solutions for addressing these issues. One other expectation that we can relate with our study is to not only contribute to the existing body of knowledge of issues, causes, and solutions in microservices systems — that lacks methodological and empirical rigor — but also to encourage other researchers to explore deeper into this highly important research area. Empirical knowledge of the nature of issues in microservices systems can help organizations to develop a better understanding of the challenges and opportunities that microservices architecture brings and how to address them.
Data availability
Link to our dataset is in the reference [19].
Acknowledgments
This work is partially sponsored by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311. The authors also thank the participants of the interviews and online survey.
Declaration of AI Assistance
During the preparation of this work, the author(s) used ChatGPT to refine grammar, improve sentence structure, and resolve formatting issues. After utilizing this tool, the author(s) thoroughly reviewed and edited the content as needed, taking full responsibility for the final publication.
[…]